### 10 Apr 2026 v3

# Fisher-Geometric Diffusion in Stochastic Gradient Descent: Optimal Rates, Oracle Complexity, and Information-Theoretic Limits

Abstract
Classical stochastic-approximation analyses treat the covariance of stochastic gradients as an exogenous modeling input. We show that under exchangeable mini-batch sampling this covariance is identified by the sampling mechanism itself: to leading order it is the projected covariance of per-sample gradients. In well-specified likelihood problems this reduces locally to projected Fisher information; for general M-estimation losses the same object is the projected gradient covariance $G^{\star}(\theta)$, which together with the Hessian induces sandwich/Godambe geometry.
This identification—not the subsequent diffusion or Lyapunov machinery, which is classical once the noise matrix is given—is the paper’s main contribution. It endogenizes the diffusion coefficient (with effective temperature $\tau=\eta/b$), determines the stationary covariance via a Lyapunov equation whose inputs are now structurally fixed, and selects the identified statistical geometry as the natural metric for convergence analysis. We prove matching upper and lower bounds of order $\Theta(1/N)$ for risk in this metric under an oracle budget $N$; the lower bound is established first via a van Trees argument in the parametric Fisher setting and then extended to adaptive oracle transcripts under a predictable-information condition and mild conditional likelihood regularity. Translating these bounds into oracle complexity yields $\varepsilon$-stationarity guarantees in the Fisher dual norm that depend on an intrinsic effective dimension $d_{\mathrm{eff}}$ and a statistical condition number $\kappa_{F}$, rather than ambient dimension or Euclidean conditioning. Numerical experiments confirm the Lyapunov predictions at both continuous-time and discrete-time levels and show that scalar temperature matching cannot reproduce directional noise structure.
Keywords: Stochastic gradient descent; Fisher information; Godambe information; information geometry; diffusion approximation; minimax rates; oracle complexity; batch-size control; simulation optimization
1 Introduction
Stochastic gradient descent (SGD) is the computational workhorse behind a large class of operations research and management science problems, including simulation optimization, large-scale stochastic programming, and data-driven estimation in service and supply-chain systems. In these settings, sampling effort is a decision variable: each iterate trades off additional replications (or scenarios) against the number of updates that can be executed under a fixed budget. Several widely observed phenomena depend on this trade-off: small mini-batches often dominate large batches in wall-clock efficiency; constant-step SGD exhibits steady-state “error floors” and implicit regularization; and plain SGD frequently behaves as if it were curvature-aware despite using only first-order information.
Classical stochastic approximation (SA) theory provides the analytical machinery—diffusion limits, Ornstein–Uhlenbeck linearization, Lyapunov stationary covariance, $\sqrt{t}$-CLT scaling—for analyzing such recursions once the gradient-noise covariance is specified. We do not claim novelty for that machinery. Our contribution is logically prior: we identify a structural property of the stochastic-gradient oracle that SA typically treats as exogenous, namely that the mini-batch gradient covariance is pinned down as $b^{-1}G^{\star}(\theta)$ under exchangeable sampling, where $G^{\star}(\theta)$ is the projected covariance of per-sample gradients. In likelihood settings this matrix reduces to Fisher information at the correctly specified point $\theta^{\star}$; for general losses it is the projected gradient covariance that, together with the Hessian $H^{\star}$, induces sandwich/Godambe geometry. Once $G^{\star}(\theta)$ is identified, the diffusion approximation for constant-step SGD is determined rather than modeled; the natural metric for rates is statistical rather than Euclidean; and batch size admits principled design rules under a fixed sampling budget. The subsequent Lyapunov analysis, rate bounds, and complexity guarantees follow by applying standard tools in this identified geometry.
1.1 Motivation: Simulation Optimization as Sampling Control
Consider the canonical simulation-optimization objective
$\min_{\theta\in\Theta}\;L(\theta)\;:=\;\mathbb{E}\big[h(\theta,\xi)\big],$ (1)
where $\xi$ denotes random inputs (arrivals, service times, demand shocks) and $h(\theta,\xi)$ is the induced cost. A standard stochastic-gradient estimator uses $b$ simulation replications,
$g_{B}(\theta)\;=\;\frac{1}{b}\sum_{i=1}^{b}\nabla_{\theta}h(\theta,\xi_{i}),\qquad\xi_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}P,$ (2)
followed by $\theta^{+}=\theta-\eta\,g_{B}(\theta)$. Classical variance-reduction intuition focuses on the scalar variance of a component or of a scalarized performance measure, suggesting that larger $b$ should always help. In practice, under fixed sampling budgets, small $b$ often wins.
The key reason is geometric: the noise in $g_{B}(\theta)$ is not scalar. Its covariance is a matrix that encodes which directions are intrinsically informative and which are weakly identified by the underlying experiment. Our results show that, to leading order, this covariance scales as $b^{-1}G^{*}(\theta)$, where $G^{*}(\theta)$ is the projected covariance of per-sample gradients. For correctly specified likelihood objectives, $G^{*}(\theta^{\star})$ reduces to projected Fisher information at the well-specified point; for general losses $G^{*}(\theta)$ is the projected gradient covariance that, together with $H^{\star}$, induces sandwich/Godambe geometry. This provides the statistical foundation for interpreting SGD as a controlled diffusion whose effective temperature is $\tau=\eta/b$.
This sampling-control perspective is not limited to simulation optimization. Three canonical MOR problem classes share the same structure: (i) simulation optimization (queuing, inventory, revenue management), where each mini-batch gradient requires $b$ simulation replications and the replication budget is the binding constraint; (ii) stochastic programming and SAA, where scenario-based gradient estimates are drawn from an empirical distribution and the number of scenarios per iterate is a design variable subject to a total scenario budget; and (iii) online M-estimation in service systems, where per-transaction loss gradients arrive in a stream and the batch window trades off estimation quality against decision latency. In each case, sampling effort is explicit, $b$ is small relative to $d$, and the question “how many samples per iterate?” is operationally binding. The theory developed below applies uniformly to all three classes.
1.2 Core Insights
Before turning to formal statements, we summarize the main idea informally. When SGD computes a mini-batch gradient by averaging $b$ per-sample gradients, the resulting noise is not generic randomness. Its covariance—the matrix describing which directions are noisy and which are quiet—is determined by the loss function and the data-generating process.
This has three immediate consequences:
  1. 1.
    Mini-batch noise is anisotropic. Doubling the batch size reduces noise equally in all directions, but the shape of the noise ellipsoid is fixed by the identified covariance $G^{\star}$, with Fisher information recovered locally under correct specification.
  2. 2.
    The natural metric for local analysis is statistical, not Euclidean. Since the noise covariance is shaped by $G^{\star}$ near $\theta^{\star}$, measuring convergence in the corresponding statistical metric gives tighter and more interpretable bounds than the Euclidean metric.
  3. 3.
    Batch size is a design variable, not just a hyperparameter. Under a fixed oracle budget $N=Tb$, batch size controls the effective temperature $\tau=\eta/b$ of the induced diffusion. It governs the stationary covariance through the Lyapunov balance, making $b$ a design lever for effort allocation.
The remainder of this section makes these claims precise.
1.3 What We Show
(i) Identification from sampling theory.
Under exchangeable sampling (and simple random mini-batching), the mini-batch gradient satisfies
$\mathrm{Cov}\!\bigl(g_{B}(\theta)\mid\mathcal{F}\bigr)\;=\;\frac{1}{b}\,G^{*}(\theta)\;+\;o(b^{-1}),$ (3)
where $G^{*}(\theta)$ is the projected gradient covariance (Section 4). For correctly specified likelihood losses, $G^{*}(\theta^{\star})=F^{\star}(\theta^{\star})$.
(ii) Diffusion and Lyapunov structure.
In the small-stepsize regime with $\tau=\eta/b$ held fixed, the SGD recursion admits a diffusion approximation whose volatility is determined by the identified covariance $G^{*}(\theta)$ (Section 5). Near a nondegenerate critical point, the limiting SDE linearizes to an OU process whose stationary covariance $\Sigma_{\infty}$ solves the Lyapunov equation
$H^{\star}\Sigma_{\infty}+\Sigma_{\infty}(H^{\star})^{\top}\;=\;\tau\,G^{*}(\theta^{\star}),$ (4)
making explicit that steady-state fluctuations are governed by the ratio of curvature to identified noise.
(iii) Matching rates and minimax lower bounds.
We analyze SGD performance in the frozen statistical geometry at $\theta^{\star}$. We establish $1/T$ upper bounds for mean-square error in the Fisher metric and match them with information-theoretic lower bounds via van Trees, showing that the rate and its dependence on intrinsic geometry are tight within the stated model class (Sections 6 and 7).
(iv) Oracle complexity in statistical conditioning.
The oracle complexity bounds scale with a statistical condition number and an effective dimension, isolating statistical difficulty from Euclidean conditioning. This clarifies when classical Hessian-based condition numbers can be misleading for statistical objectives.
(v) Batch size as an OR design variable.
The framework above connects batch-size design to the identified statistical geometry; Section 9 discusses extensions to misspecification, dependence, and operational effort allocation.
Main Results (informal).
Theorem 4.3 — Identification of mini-batch covariance under exchangeable sampling: $\mathrm{Cov}(g_{B})=b^{-1}G^{\star}(\theta)+o(b^{-1})$. Theorem 5.4 — Diffusion approximation with volatility determined by $G^{\star}$. Thm. 5.6, Cor. 5.7 — OU linearization at $\theta^{\star}$; stationary covariance solves a Lyapunov equation. Thm. 6.5, Prop. 6.10 — Matching $1/(Tb)$ upper and lower bounds in the frozen Fisher metric. Theorem 7.3 — Oracle complexity $N\!=\!\widetilde{\Theta}(\kappa_{F}\,d_{\mathrm{eff}}/\varepsilon^{2})$ for Fisher-dual stationarity.
1.4 Organization
Section 2 reviews related work. Section 3 defines the oracle model and geometric objects. Section 4 derives the covariance identification from exchangeability. Section 5 builds the diffusion approximation, OU linearization, and discrete-time bridge. Section 6 proves matching upper and lower bounds in Fisher geometry, and Section 7 states oracle complexity. Section 8 provides numerical validation, and Section 9 discusses extensions and practical implications.
2 Literature Review and Positioning
We organize prior work along two axes: (i) whether the mini-batch noise covariance is assumed (exogenous) or derived from the sampling mechanism; and (ii) whether rates are stated in a Euclidean or in the identified statistical metric. Table 1 summarizes the resulting taxonomy.
Stochastic approximation and stationary covariance.
Classical SA theory studies recursions $\theta_{t+1}=\theta_{t}-\eta_{t}\{h(\theta_{t})+\xi_{t+1}\}$ and derives a $\sqrt{t}$-CLT whose asymptotic covariance is governed by a Lyapunov equation involving the Jacobian and a long-run noise covariance $\Gamma$ (Robbins and Monro, 1951; Dvoretzky, 1956; Kushner and Yin, 2003; Polyak and Juditsky, 1992). This framework takes $\Gamma$ as given; the OU linearization, Lyapunov balance, and $1/\sqrt{t}$ scaling are standard consequences. Our contribution is logically prior: we derive $\Gamma=G^{\star}(\theta)/b$ from the mini-batch sampling design (Theorem 4.3), converting the SA machinery from a conditional analysis into a structural one.
Diffusion approximations for SGD.
A substantial literature interprets constant-step SGD as a discretization of an SDE with temperature $\tau=\eta/b$ and various proposed diffusion matrices $\mathbf{C}(\theta)$ (Mandt et al., 2017; Li et al., 2017; Smith et al., 2021). In those works the diffusion covariance is either postulated, estimated from running iterates, or justified heuristically through moment matching. Our covariance identification removes this modeling degree of freedom: under exchangeable mini-batching, the diffusion matrix is determined as $G^{\star}(\theta)$ (Theorem 5.4), and the local OU regime inherits its Lyapunov balance at temperature $\tau$ (Corollary 5.7). Because the identification is logically prior to the discretization corrections of Li et al. (2017) or the Hessian-scaling regime of Smith et al. (2021), it is what enables the subsequent rate-matching bounds.
Scaling regimes.
Recent rigorous diffusion-limit work studies joint regimes where $b$ grows with $d$ (Paquette et al., 2024). Our analysis is complementary and closer to OR operating constraints: we keep $d$ fixed and treat $b$ as a design variable subject to a budget, with particular attention to the common regime $b\ll d$ where the matrix geometry of $G^{\star}(\theta)$ controls transient behavior and risk.
Information geometry, natural gradients, and Fisher-shaped preconditioning.
Information geometry treats Fisher information as a Riemannian metric and motivates natural-gradient updates $\theta_{t+1}=\theta_{t}-\eta\,F(\theta_{t})^{-1}\nabla L(\theta_{t})$ (Amari, 1998; Martens, 2020). Related methods—including stochastic gradient Fisher scoring (Ahn et al., 2012) and KFAC (Martens and Grosse, 2015)—inject or exploit Fisher geometry exogenously in the drift or preconditioner. Our contribution is distinct: even plain SGD, without any preconditioning, inherits the identified statistical geometry endogenously through its noise covariance under mini-batch sampling, changing the interpretation of diffusion strength, stationary risk, and batch-size tuning without requiring natural-gradient updates.
Minimax rates and information-theoretic limits.
Classical oracle lower bounds are stated in Euclidean norms under smoothness and strong convexity (Nemirovski et al., 2009; Agarwal et al., 2012), which may not reflect the intrinsic difficulty of statistical objectives. We state both upper and lower bounds in the identified statistical metric via a van Trees argument (Theorem 6.5 and Proposition 6.10), with complexity governed by statistical curvature and $d_{\operatorname{eff}}$ rather than ambient dimension.
OR context: sampling budgets.
In simulation optimization and stochastic programming, gradients are estimated from sampled scenarios and sampling cost is explicit (Fu, 2006; Shapiro et al., 2014; Chernoff, 1959). Because batching controls temperature $\tau=\eta/b$, it directly shapes diffusion amplitude and terminal risk; Section 9 outlines the resulting effort allocation problem.
Why small batches can work when $b\ll d$.
The empirical preference for small batches (Keskar et al., 2017; Shallue et al., 2019; McCandlish et al., 2018) is explained by our framework: smaller $b$ buys more contraction steps under fixed budgets, the identified covariance concentrates noise along statistically flat directions, and rates scale with $d_{\operatorname{eff}}\ll d$.
Table 1: Positioning: primitives, contrasts, and where each appears in the paper.
Primitive Status quo This paper Pointers
Mini-batch noise covariance Exogenous / isotropic / heuristic $\Gamma$ Identified from sampling: $\mathrm{Cov}(g_{B})\propto b^{-1}G^{\star}(\theta)$; Fisher is a special case Sec. 4; Thm. 4.3
Diffusion matrix for SGD Postulated $\mathbf{C}(\theta)$ in SDEs Diffusion coefficient determined by sampling theory: $G^{\star}(\theta)$ Sec. 5; Thm. 5.4
OU / Lyapunov equilibrium Heuristic OU with ad hoc covariance OU linearization with Lyapunov balance driven by $G^{\star}$ at temperature $\tau=\eta/b$ Sec. 5; Cor. 5.7
Metric for rates Euclidean distance; $\kappa_{H}$ dominates Identified statistical metric is the natural risk Sec. 6; Thm. 6.5
Lower bounds Euclidean oracle lower bounds Information-theoretic lower bound in Fisher metric (van Trees) Sec. 6; Prop. 6.10
Dimension notion Ambient $d$ Effective dimension $d_{\operatorname{eff}}$ induced by $F^{\star}$ / $G^{\star}$ Sec. 6; Def. (36)
Oracle complexity Euclidean stationarity; $\kappa_{H}$ Fisher-dual stationarity; $\kappa_{F}$ and $d_{\operatorname{eff}}$ Sec. 7; Thm. 7.3
3 Problem Setup and Preliminaries
This section fixes notation, specifies the stochastic-gradient oracle, and records the geometric objects that govern the diffusion approximation and the rate/complexity bounds. The perspective is Mathematics of Operations Research: we make the sampling design explicit, separate intrinsic noise geometry from finite-population correction, and state the local regularity conditions used throughout.
3.1 Notation and macros used throughout
We work on a measurable space $(\mathcal{Z},\mathcal{A})$. Vectors use the Euclidean inner product $\langle u,v\rangle=u^{\top}v$ and norm $\|u\|_{2}$. For a symmetric matrix $A$, $\|A\|_{\operatorname{op}}$ denotes the operator norm and $\lambda_{\min}(A),\lambda_{\max}(A)$ its extreme eigenvalues.
3.2 Data model, objective, and stochastic-gradient oracle
Let $\Theta\subseteq\mathbb{R}^{d}$ be open and convex. Let $Q$ denote the (unknown) data distribution on $(\mathcal{Z},\mathcal{A})$, and let $\ell:\Theta\times\mathcal{Z}\to\mathbb{R}$ be a measurable loss. Define the population risk
$L(\theta)\;:=\;\mathbb{E}_{Z\sim Q}\bigl[\ell(\theta;Z)\bigr],$ (5)
and let $\theta^{\star}\in\arg\min_{\theta\in\Theta}L(\theta)$ be a (local) minimizer.
Finite-population sampling model.
We observe a dataset $z_{1:n}:=(z_{1},\dots,z_{n})$ drawn i.i.d. from $Q$ (or, in the exchangeability formulation of Section 4, conditionally i.i.d. given a directing measure). All lower bounds in Section 6 are stated for a general oracle transcript with bounded predictable information (Assumption 6.9); the i.i.d. sampling model is recovered as a special case. At iteration $t$, the algorithm selects a mini-batch $B_{t}\subseteq\{1,\dots,n\}$ of size $b_{t}$ uniformly without replacement and forms the mini-batch gradient
$g_{t}(\theta)\;:=\;\frac{1}{b_{t}}\sum_{i\in B_{t}}\psi(\theta;z_{i}),\qquad\psi(\theta;z):=\nabla_{\theta}\ell(\theta;z).$ (6)
The SGD recursion is
$\theta_{t+1}\;=\;\theta_{t}-\eta_{t}\,g_{t}(\theta_{t}),$ (7)
with stepsize $\eta_{t}>0$.
Filtration and martingale structure.
Let $\mathcal{F}_{t}$ be the $\sigma$-field generated by $(\theta_{0},B_{0},\dots,B_{t-1},\theta_{t})$ and the dataset $z_{1:n}$. Then
$\xi_{t}(\theta_{t}):=g_{t}(\theta_{t})-\mathbb{E}[g_{t}(\theta_{t})\mid\mathcal{F}_{t}]$
is a martingale difference sequence with respect to $(\mathcal{F}_{t})$, and all diffusion limits are driven by the predictable quadratic variation of $\sum_{t}\eta_{t}\xi_{t}$.
Cost accounting (oracle calls).
We measure computational cost in sample-gradient evaluations. One iteration with batch size $b_{t}$ uses $b_{t}$ sample evaluations. After $T$ iterations, the total sampling budget is
$N_{T}\;:=\;\sum_{t=0}^{T-1}b_{t}.$ (8)
We report results both in iterations $T$ and in oracle calls $N_{T}$, since control/OR questions naturally optimize over batch allocation under a fixed $N_{T}$.
3.3 Likelihood losses: Fisher information and the induced metric
When $\ell(\theta;z)=-\log p_{\theta}(z)$ for a parametric family $\{P_{\theta}:\theta\in\Theta\}$ with density $p_{\theta}$, the score is $s_{\theta}(z):=\nabla_{\theta}\log p_{\theta}(z)$ and the loss gradient is the negative score, $\nabla_{\theta}\ell(\theta;z)=-s_{\theta}(z)$. The intrinsic noise geometry reduces to Fisher information at the correctly specified point $\theta^{\star}$.
Definition 3.1 (Fisher information).
Assume $p_{\theta}$ is differentiable in $\theta$ and $s_{\theta}(Z)$ is square integrable under $P_{\theta}$. The Fisher information is
$F(\theta)\;:=\;\mathbb{E}_{\theta}\!\bigl[s_{\theta}(Z)s_{\theta}(Z)^{\top}\bigr].$ (9)
Lemma 3.2 (Score identities (Lehmann and Casella, 1998, Ch. 7)).
Assume $\int p_{\theta}(z)\,dz=1$ and differentiation under the integral is valid. Then (i) $\mathbb{E}_{\theta}[s_{\theta}(Z)]=0$ and (ii)
$F(\theta)\;=\;-\mathbb{E}_{\theta}\!\bigl[\nabla_{\theta}^{2}\log p_{\theta}(Z)\bigr].$ (10)
Proof.
Standard; see, e.g., Lehmann and Casella (1998, Ch. 7). ∎
Fisher metric.
If $F(\theta)\succ 0$, it defines the local inner product $\langle u,v\rangle_{F(\theta)}:=u^{\top}F(\theta)v$ and norm $\|u\|_{F(\theta)}:=\sqrt{u^{\top}F(\theta)u}$. On compact sets where $F(\theta)$ is bounded and bounded away from 0, Fisher and Euclidean norms are equivalent.
3.4 General losses: sandwich / Godambe geometry
For a general loss, information equality need not hold. The relevant local geometry splits into (i) curvature of the drift and (ii) covariance of the estimating equation (gradient).
Definition 3.3 (Noise covariance, sensitivity, and Godambe information).
Let $\psi(\theta;z):=\nabla_{\theta}\ell(\theta;z)$ and define
$G(\theta)$ $:=\mathrm{Cov}\bigl(\psi(\theta;Z)\bigr),$ (11)
$H(\theta)$ $:=\nabla^{2}L(\theta)\;=\;\mathbb{E}\!\left[\nabla_{\theta}\psi(\theta;Z)\right],$ (12)
where expectations are under $Z\sim Q$. When $H(\theta)$ is invertible on the identifiable directions, the classical sandwich covariance and Godambe information are
$\Sigma_{\mathrm{sand}}(\theta):=H(\theta)^{-1}G(\theta)H(\theta)^{-1},\qquad J_{\mathrm{God}}(\theta):=\Sigma_{\mathrm{sand}}(\theta)^{-1}=H(\theta)^{\top}G(\theta)^{-1}H(\theta).$ (13)
Remark 3.4 (Likelihood as a special case).
If $\ell(\theta;z)=-\log p_{\theta}(z)$ and the model is correctly specified, then $\psi=s_{\theta}$ and $G(\theta)=H(\theta)=F(\theta)$ at $\theta^{\star}$, so the sandwich collapses and the Godambe geometry equals Fisher.
3.5 Projection onto identifiable directions
Projection is only needed when the parametrization is redundant or when only a lower-dimensional tangent space is identifiable.
Definition 3.5 (Tangent-space projection and projected geometry).
Let $T_{\theta}\subseteq\mathbb{R}^{d}$ denote the identifiable tangent space at $\theta$ (i.e., the linear span of directions along which the loss varies at first order; formally $T_{\theta}:=\mathrm{range}(H(\theta))$, or equivalently $\mathrm{supp}(F(\theta))$ in the likelihood case), and let $\Pi_{\theta}$ be the Euclidean-orthogonal projector onto $T_{\theta}$. Define the projected noise covariance
$G^{\star}(\theta)\;:=\;\Pi_{\theta}\,G(\theta)\,\Pi_{\theta}^{\top},$ (14)
and, in the likelihood case, the projected Fisher information
$F^{\star}(\theta)\;:=\;\Pi_{\theta}\,F(\theta)\,\Pi_{\theta}^{\top}.$ (15)
When $T_{\theta}=\mathbb{R}^{d}$, $\Pi_{\theta}=I$ and $G^{\star}=G$, $F^{\star}=F$.
Remark 3.6 (Terminological convention: “Fisher/Godambe”).
Throughout the paper we write “Fisher/Godambe” as shorthand for the following identification: in the correctly specified likelihood case, $G^{\star}(\theta^{\star})=F^{\star}(\theta^{\star})$ (Fisher information); for general losses, $G^{\star}(\theta^{\star})$ is the projected gradient covariance whose inverse is the Godambe information. Every theorem statement holds with $G^{\star}$ in the general case and specializes to $F^{\star}$ under correct specification. When we write “Fisher metric” or “Fisher-dual norm,” the general-loss reader should substitute the corresponding Godambe/sandwich object.
Remark 3.7 (What enters later sections).
Sections 45 show that the mini-batch noise covariance is (up to a sampling-design factor) exactly $G^{\star}(\theta)/b_{t}$ and reduces to $F^{\star}(\theta)/b_{t}$ under correct specification. This is the only matrix that enters the diffusion coefficient and Lyapunov balance.
3.6 Local regularity conditions
All results are local near $\theta^{\star}$ and use standard smoothness/moment and nondegeneracy conditions.
Assumption 3.8 (Smoothness and moments).
The loss $\ell(\theta;z)$ is twice continuously differentiable in $\theta$. There exist constants $M<\infty$ and $m_{4}<\infty$ such that for all $\theta$ in a neighborhood $\mathcal{N}$ of $\theta^{\star}$:
  1. (a)
    $\|\nabla_{\theta}^{2}\ell(\theta;z)\|_{\operatorname{op}}\leq M$ almost surely;
  2. (b)
    $\mathbb{E}\bigl[\|\psi(\theta;Z)\|_{2}^{4}\bigr]\leq m_{4}$.
Assumption 3.9 (Local stability).
The population risk is locally strongly convex at $\theta^{\star}$:
$H^{\star}:=\nabla^{2}L(\theta^{\star})\succeq\mu I\quad\text{for some }\mu>0.$ (16)
Assumption 3.10 (Geometric nondegeneracy).
The intrinsic noise geometry is positive definite on identifiable directions:
$G^{\star}(\theta^{\star})\succeq\lambda_{\min}I\quad\text{for some }\lambda_{\min}>0.$ (17)
(In the correctly specified likelihood case this is $F^{\star}(\theta^{\star})\succ 0$.)
Remark 3.11 (Role of assumptions).
Assumptions 3.83.10 provide Taylor/moment control, a locally stable equilibrium, and nondegeneracy of the diffusion coefficient, respectively.
3.7 Key geometric parameters (preview)
Several quantities derived from the local geometry at $\theta^{\star}$ govern the rate and complexity results of Sections 67. We collect their definitions here so the reader can refer to them before encountering the formal theorems.
Definition 3.12 (Fisher-strong-convexity constant).
Let $H^{\star}:=\nabla^{2}L(\theta^{\star})$ and $F^{\star}:=F(\theta^{\star})$ (or $G^{\star}(\theta^{\star})$ for general losses). Define
$\mu_{F}\;:=\;\lambda_{\min}\!\left(\frac{(F^{\star})^{1/2}H^{\star}(F^{\star})^{-1/2}+(F^{\star})^{-1/2}H^{\star}(F^{\star})^{1/2}}{2}\right).$ (18)
Equivalently, $\mu_{F}$ is the smallest eigenvalue of the symmetrized curvature–information interaction $\mathrm{Sym}((F^{\star})^{1/2}H^{\star}(F^{\star})^{-1/2})$.
Definition 3.13 (Fisher condition number and effective dimension).
$\kappa_{F}\;:=\;\frac{\lambda_{\max}(F^{\star})}{\lambda_{\min}(F^{\star})},\qquad d_{\mathrm{eff}}(F^{\star})\;:=\;\frac{\mathrm{Tr}(F^{\star})}{\lambda_{\max}(F^{\star})}\;\in\;(0,d].$ (19)
The quantity $\kappa_{F}$ measures statistical anisotropy rather than Euclidean ill-conditioning; a problem can be Euclidean-stiff ($\kappa_{H}\gg 1$) yet statistically well-conditioned ($\kappa_{F}=O(1)$) when curvature and information share eigenstructure. The effective dimension $d_{\mathrm{eff}}$ is the stable rank of $F^{\star}$: it equals $d$ when $F^{\star}$ is isotropic and approaches $1$ when a single direction dominates.
Remark 3.14 (Generalized eigenvalue interpretation).
The constant $\mu_{F}$ is linked to the generalized eigenvalues of the pair $(H^{\star},F^{\star})$. If $H^{\star}$ and $F^{\star}$ commute (i.e., share an eigenbasis $\{v_{j}\}$ with respective eigenvalues $h_{j}$ and $f_{j}$), then $(F^{\star})^{1/2}H^{\star}(F^{\star})^{-1/2}$ has eigenvalues $f_{j}^{1/2}\,h_{j}\,f_{j}^{-1/2}=h_{j}$, so the symmetrization is trivial and $\mu_{F}=\min_{j}h_{j}=\lambda_{\min}(H^{\star})$: the Fisher-strong-convexity constant reduces to the ordinary strong-convexity constant. In the non-commuting case, $(F^{\star})^{1/2}H^{\star}(F^{\star})^{-1/2}$ is non-normal (its eigenvalues are still those of $H^{\star}$, but it is no longer symmetric), and the minimum of its quadratic form—i.e., $\mu_{F}$—can drop strictly below $\lambda_{\min}(H^{\star})$. Geometrically, this occurs when the Fisher metric amplifies directions along which curvature is weak, making the effective contraction rate slower than the Euclidean strong-convexity constant. The product $\kappa_{F}\cdot d_{\mathrm{eff}}$ that enters the oracle complexity (Theorem 7.3) therefore measures the total cost of resolving all statistically relevant directions at the hardest direction’s scale.
Table 2: Key notation. All quantities are evaluated locally at or near $\theta^{\star}$ unless noted otherwise.
Symbol Meaning
$L(\theta)=\mathbb{E}[\ell(\theta;Z)]$ Population risk (loss)
$\psi(\theta;z)=\nabla_{\theta}\ell(\theta;z)$ Per-sample estimating equation (gradient)
$g_{B}(\theta)$ Mini-batch gradient estimator
$G(\theta)=\mathrm{Cov}(\psi(\theta;Z))$ Gradient covariance; $G^{\star}=\Pi G\Pi^{\top}$ projected
$F(\theta)=\mathbb{E}[s_{\theta}s_{\theta}^{\top}]$ Fisher information (likelihood case)
$F^{\star}(\theta^{\star})$ Projected Fisher; equals $G^{\star}(\theta^{\star})$ under correct specification
$H^{\star}=\nabla^{2}L(\theta^{\star})$ Hessian at optimum (drift curvature)
$J_{\mathrm{God}}=H^{\top}G^{-1}H$ Godambe (sandwich) information
$\tau=\eta/b$ Effective temperature
$\mu_{F}$ Fisher-strong-convexity constant (Def. 3.12)
$\kappa_{F}$ Fisher condition number (Def. 3.13)
$d_{\operatorname{eff}}(F^{\star})$ Effective dimension / stable rank of $F^{\star}$
4 Fisher/Godambe Alignment of Mini-Batch Noise
Mini-batch gradient noise is not an arbitrary covariance: it is determined by (i) the sampling design used to form mini-batches and (ii) the intrinsic covariance of the per-sample estimating equation $\psi(\theta;Z)$ (Fisher information in the well-specified likelihood case, the projected gradient covariance more generally). We make this precise in two steps:
  • an exact finite-population identity conditional on the realized dataset (classical survey-sampling algebra), and
  • an exchangeable/i.i.d. limit that converts the dataset-dependent quantity into the population covariance $G(\theta):=\mathrm{Cov}(\psi(\theta;Z))$ (or its projected analogue).
4.1 Finite-population covariance identity (exact, conditional on data)
Fix $\theta$ and condition on the realized dataset $\{z_{1},\dots,z_{n}\}$. Write $\psi_{i}:=\psi(\theta;z_{i})$ and define the finite-population mean and covariance
$\bar{\psi}_{n}(\theta):=\frac{1}{n}\sum_{i=1}^{n}\psi_{i},\qquad S_{n}(\theta):=\frac{1}{n-1}\sum_{i=1}^{n}(\psi_{i}-\bar{\psi}_{n})(\psi_{i}-\bar{\psi}_{n})^{\top}.$
Let $B$ be a simple random sample without replacement of size $b$ from $\{1,\dots,n\}$ and define the mini-batch average
$g_{B}(\theta):=\frac{1}{b}\sum_{i\in B}\psi_{i}.$
Lemma 4.1 (Exact finite-population covariance of the mini-batch mean).
Conditionally on the dataset,
$\mathrm{Cov}\!\bigl(g_{B}(\theta)\mid z_{1:n}\bigr)\;=\;\frac{1}{b}\Bigl(1-\frac{b}{n}\Bigr)\,S_{n}(\theta).$ (20)
Proof.
Write $g_{B}(\theta)=\frac{1}{b}\sum_{i=1}^{n}I_{i}\,\psi_{i}$ where $I_{i}=\mathbf{1}\{i\in B\}$. Under simple random sampling without replacement, $\mathbb{E}[I_{i}]=b/n$, $\mathrm{Var}(I_{i})=\frac{b}{n}(1-\frac{b}{n})$, and for $i\neq j$, $\mathrm{Cov}(I_{i},I_{j})=-\frac{b(n-b)}{n^{2}(n-1)}$. Expanding $\mathrm{Cov}(\sum_{i}I_{i}\psi_{i}\mid z_{1:n})$ and collecting terms yields (20). ∎
Remark 4.2 (Equivalent correction-factor forms).
Using $S_{n}$ with the 1/(n1)1/(n-1) normalization gives the factor (1b/n)(1-b/n). If instead one uses the $1/n$ covariance normalization, the factor becomes 1(b1)/(n1)1-(b-1)/(n-1). The two are algebraically equivalent up to the change of normalization.
Lemma 4.1 is the clean MOR-style decomposition: sampling design contributes the scalar finite-population correction 1b/n1-b/n, while the geometry lives in $S_{n}(\theta)$.
4.2 From exchangeability to Godambe/Fisher alignment (population geometry)
We now convert the dataset-dependent covariance $S_{n}(\theta)$ into a population-level object.
Let $(Z_{i})_{i\geq 1}$ be exchangeable with de Finetti directing measure $\mu$ (Hewitt and Savage, 1955). Conditional on $\mu$, the sequence is i.i.d. with law $P_{\mu}$. Define the population covariance (conditional on $\mu$)
$G_{\mu}(\theta)\;:=\;\mathrm{Cov}_{\mu}\bigl(\psi(\theta;Z)\bigr),\qquad Z\sim P_{\mu},$ (21)
and, when identifiability requires it, the projected version $G_{\mu}^{\star}(\theta):=\Pi_{\theta}G_{\mu}(\theta)\Pi_{\theta}^{\top}$.
Theorem 4.3 (Godambe alignment under exchangeability).
Assume $\mathbb{E}_{\mu}\|\psi(\theta;Z)\|_{2}^{2}<\infty$. Let $g_{B}(\theta)=\frac{1}{b}\sum_{i\in B}\psi(\theta;Z_{i})$ where $B$ is a uniform subset of $\{1,\dots,n\}$ of size $b$ (sampling without replacement). Then, conditional on $\mu$,
$\mathrm{Cov}\!\bigl(g_{B}(\theta)\mid\mu\bigr)\;=\;\frac{1}{b}\,G_{\mu}(\theta).$ (22)
Interpretation. Equation (22) is exact (not an approximation) when one marginalizes over both the mini-batch index $B$ and the dataset realization conditional on $\mu$. The finite-population correction (1b/n)(1-b/n) of Lemma 4.1 is present when one conditions on a fixed dataset, but it disappears at the population level because the randomness of the dataset itself contributes a complementary term through the total-covariance decomposition. Equation (23) below makes this bookkeeping explicit.
Moreover,
$\mathbb{E}\!\Big[\mathrm{Cov}(g_{B}(\theta)\mid Z_{1:n},\mu)\,\big|\,\mu\Big]\;=\;\frac{1}{b}\Bigl(1-\frac{b}{n}\Bigr)\,\mathbb{E}[S_{n}(\theta)\mid\mu]\;=\;\frac{1}{b}\Bigl(1-\frac{b}{n}\Bigr)\,G_{\mu}(\theta),$ (23)
and the remaining part of (22) comes from the randomness of the dataset through the total-covariance decomposition.
Proof.
Conditional on $\mu$, the sequence $(Z_{i})_{i\geq 1}$ is i.i.d. with law $P_{\mu}$; therefore any uniformly sampled subset of size $b$ has the same joint distribution as the first $b$ coordinates. Hence selecting indices $B$ without replacement from $\{1,\dots,n\}$ is distributionally equivalent to taking the first $b$ terms, and $g_{B}(\theta)$ is the average of $b$ i.i.d. draws from $P_{\mu}$. This yields $\mathrm{Cov}(g_{B}(\theta)\mid\mu)=\frac{1}{b}G_{\mu}(\theta)$, proving (22).
For (23), apply the exact identity Lemma 4.1 conditional on $(Z_{1:n},\mu)$ and take $\mathbb{E}[\cdot\mid\mu]$. Under $\mathbb{E}_{\mu}\|\psi\|^{2}<\infty$, $\mathbb{E}[S_{n}(\theta)\mid\mu]=G_{\mu}(\theta)$ (the 1/(n1)1/(n-1) sample-covariance normalization is unbiased). ∎
Remark 4.4 (What “alignment” means operationally).
Equation (22) is the alignment statement used by the diffusion/Lyapunov analysis: mini-batch noise has covariance $\frac{1}{b}$ times a population matrix determined by the loss and the data law. The finite-population correction in Lemma 4.1 matters when one conditions on a fixed dataset and reuses it many times; in the standard stochastic-approximation regime with fresh randomness across iterations, the population covariance (22) is the relevant object.
Proposition 4.5 (Finite-population correction under dataset reuse).
Under the notation of Theorem 4.3, three conditioning levels yield distinct covariance expressions:
  1. (i)
    Conditional on the dataset $z_{1:n}$ (fixed pool, sampling without replacement): $\mathrm{Cov}(g_{B}(\theta)\mid z_{1:n})=\frac{1}{b}(1-\frac{b}{n})\,S_{n}(\theta)$ (Lemma 4.1).
  2. (ii)
    Conditional on the directing measure $\mu$ (fresh data per iteration): $\mathrm{Cov}(g_{B}(\theta)\mid\mu)=\frac{1}{b}\,G_{\mu}(\theta)$ (Theorem 4.3), with no finite-population correction.
  3. (iii)
    Population level (unconditional): $\mathrm{Cov}(g_{B}(\theta))=\frac{1}{b}\,\mathbb{E}_{\mu}[G_{\mu}(\theta)]+\mathrm{Cov}_{\mu}(\mathbb{E}[g_{B}(\theta)\mid\mu])$. This reduces to $\frac{1}{b}\,G(\theta)$ when the conditional mean $\mathbb{E}[\psi(\theta;Z)\mid\mu]$ does not vary with $\mu$ (e.g., under identifiability with a single data-generating distribution).
When a finite dataset is reused across many iterations (i.e., without-replacement sampling from the same $z_{1:n}$ at every step), the conditional covariance carries the factor (1b/n)(1-b/n), which is non-negligible unless $b=o(n)$. The main diffusion, rate, and complexity results (Sections 57) are population-level or fresh-sampling statements that use level (ii).
Proof.
Level (i) is Lemma 4.1. Level (ii) is Theorem 4.3. Level (iii) follows by integrating level (ii) over $\mu$ under identifiability. The total-covariance decomposition $\mathrm{Cov}(g_{B})=\mathbb{E}_{\mu}[\mathrm{Cov}(g_{B}\mid\mu)]+\mathrm{Cov}_{\mu}(\mathbb{E}[g_{B}\mid\mu])$ reconciles levels (ii) and (iii), with the second term contributing when $\mathbb{E}[g_{B}\mid\mu]$ varies with $\mu$. ∎
Remark 4.6 (Batch size rescales temperature, not geometry).
Equation (22) implies that increasing the batch size $b$ rescales the effective temperature $\tau=\eta/b$ (Section 5) but does not alter the noise ellipsoid: its shape is entirely determined by the statistical geometry $G_{\mu}(\theta)$. This decoupling—scale from the sampling budget, shape from the loss—underlies the subsequent rate and complexity results.
[…]
Figure 1: Isotropic noise (left) vs. Fisher-aligned noise (right) at the same total variance. Batch size rescales temperature but not shape.
Remark 4.7 (Specialization to Fisher alignment).
If $\ell(\theta;z)=-\log p_{\theta}(z)$ and the model is correctly specified, then $\psi(\theta;z)=s_{\theta}(z)$ is the score and the Bartlett identity yields $\mathrm{Cov}(s_{\theta}(Z))=F(\theta)$ at $\theta=\theta^{\star}$. Projecting onto the identifiable tangent space gives $G_{\mu}^{\star}(\theta^{\star})=F^{\star}(\theta^{\star})$. Thus (22) becomes $\mathrm{Cov}(g_{B}(\theta^{\star})\mid\mu)=\frac{1}{b}F^{\star}(\theta^{\star})$, i.e., Fisher alignment.
Remark 4.8 (No Gaussian assumption).
Neither Lemma 4.1 nor Theorem 4.3 assumes Gaussian noise. Gaussianity enters later only as a diffusion approximation (martingale CLT) when we pass from the discrete recursion to an OU limit.
5 Continuous-Time Diffusion Approximation and the OU Regime
This section links the discrete-time SGD recursion to a continuous-time SDE whose diffusion covariance is determined by the mini-batch sampling model and the intrinsic noise geometry derived in Section 4. Specifically, $G^{\star}(\theta)$ was defined in Section 3 and shown in Theorem 4.3 to govern the leading-order mini-batch covariance. Near a nondegenerate critical point, the SDE linearizes to an Ornstein–Uhlenbeck (OU) process, and its stationary covariance is characterized by a Lyapunov equation. The resulting “temperature” parameter $\tau=\eta/b$ is the scalar that converts sampling effort into diffusion strength. Diffusion approximations are a standard tool in OR for reducing complex stochastic systems to tractable Brownian models (Harrison, 1985; Whitt, 2002); the present setting is analogous, with the sampling mechanism playing the role of the arrival process.
The stationary covariance and risk expressions derived below depend only on the local second-moment structure of the oracle noise and do not require independence across oracle calls; a martingale-difference oracle (Assumption 6.9) suffices.
5.1 SGD decomposition and effective temperature
Fix constant stepsize $\eta>0$ and batch size $b$. Let
$h(\theta):=\nabla L(\theta)=\mathbb{E}[\psi(\theta;Z)],\qquad\psi(\theta;z):=\nabla_{\theta}\ell(\theta;z),$
and write the mini-batch gradient estimator as
$g_{t}(\theta_{t})=h(\theta_{t})+\xi_{t}(\theta_{t}),\qquad\mathbb{E}[\xi_{t}(\theta_{t})\mid\mathcal{F}_{t}]=0,$
where $\mathcal{F}_{t}$ is the natural filtration. Then SGD is
$\theta_{t+1}=\theta_{t}-\eta\,h(\theta_{t})-\eta\,\xi_{t}(\theta_{t}).$ (24)
By Theorem 4.3 (and its Fisher specialization in Remark 4.7), the conditional covariance of the noise obeys
$\mathrm{Cov}\!\bigl(\xi_{t}(\theta)\mid\mathcal{F}_{t}\bigr)=\frac{1}{b}\,G^{\star}(\theta)+r_{b,n}(\theta),\qquad\|r_{b,n}(\theta)\|_{\operatorname{op}}=o(b^{-1})+O(n^{-1}),$ (25)
uniformly in a local neighborhood (the $O(n^{-1})$ term accounts for finite-population corrections and covariance-estimation bias; it is negligible whenever $b=o(n)$).
We call $\tau:=\eta/b$ the effective temperature:
$\tau:=\eta/b,$ (26)
which will appear as the single scalar factor multiplying the diffusion covariance in the continuous-time limit.
5.2 A diffusion approximation (Euler–Maruyama viewpoint)
Let $s=t\eta$ denote algorithmic time and define the piecewise-linear interpolation $\bar{\theta}^{\eta}(\cdot)$ by $\bar{\theta}^{\eta}(t\eta)=\theta_{t}$ and linearity on $t\eta,(t+1)\eta]$.
Assumption 5.1 (Local regularity and localization).
On a neighborhood $\mathcal{U}$ of $\theta^{\star}$:
  1. (i)
    $h(\theta):=\nabla L(\theta)$ is locally Lipschitz;
  2. (ii)
    $\nabla^{2}L(\theta)$ is locally Lipschitz (i.e., there exists $L_{H}<\infty$ such that $\|\nabla^{2}L(\theta)-\nabla^{2}L(\theta^{\prime})\|_{\operatorname{op}}\leq L_{H}\|\theta-\theta^{\prime}\|_{2}$ for all $\theta,\theta^{\prime}\in\mathcal{U}$);
  3. (iii)
    $\sup_{\theta\in\mathcal{U}}\mathbb{E}\|\psi(\theta;Z)\|^{4}<\infty$ (bounded fourth moment; implied by Assumption 3.8(b));
  4. (iv)
    $H^{\star}:=\nabla^{2}L(\theta^{\star})\succ 0$ (local strong convexity; Assumption 3.9);
  5. (v)
    for each fixed horizon $T>0$, $\Pr(\theta_{t}\in\mathcal{U}\ \forall\,t\leq T/\eta)\to 1$ as $\eta\downarrow 0$ (localization; standard in stochastic approximation).
Assumption 5.2 (Asymptotic conditional covariance and Lindeberg).
Uniformly on $\mathcal{U}$, $\mathrm{Cov}(\xi_{t}(\theta)\mid\mathcal{F}_{t})=b^{-1}G^{\star}(\theta)+o(b^{-1})$ and the conditional Lindeberg condition holds for the martingale increments $\eta\xi_{t}(\theta_{t})$ (e.g., via a conditional CLT under exchangeability as in Section 4).
Remark 5.3 (Verifiable sufficient conditions).
Assumption 5.2 holds if $\sup_{\theta\in\mathcal{U}}\mathbb{E}[\|\psi(\theta;Z)\|_{2}^{2+\delta}]<\infty$ for some $\delta>0$ (guaranteed by Assumption 3.8(b)), $G^{\star}(\theta)$ is continuous on $\mathcal{U}$, and the mini-batch is drawn by simple random sampling from an exchangeable sequence.
Theorem 5.4 (Diffusion approximation).
Suppose Assumptions 5.15.2 hold. For fixed batch size $b\geq 1$ and constant step size $\eta>0$, define the effective temperature $\tau:=\eta/b$. For sufficiently small $\eta$, the SGD recursion admits a diffusion approximation: the piecewise-linear interpolation $\bar{\theta}^{\eta}(\cdot)$ converges weakly on $D([0,T];\mathbb{R}^{d})$ as $\eta\downarrow 0$ to the solution $\theta(\cdot)$ of the SDE
$d\theta_{s}=-h(\theta_{s})\,ds+\sqrt{\tau}\,C^{\star}(\theta_{s})\,dW_{s},\qquad C^{\star}(\theta)\,{C^{\star}(\theta)}^{\top}=G^{\star}(\theta),$ (27)
where $C^{\star}$ is, e.g., the symmetric PSD square root of $G^{\star}$,
where $W_{s}$ is standard $d$-dimensional Brownian motion. In particular, the diffusion has instantaneous covariance $\tau\,G^{\star}(\theta_{s})\,ds$. In the correctly specified likelihood case, $G^{\star}(\theta^{\star})=F^{\star}(\theta^{\star})$ at $\theta^{\star}$, so the diffusion covariance is locally Fisher-structured near the optimum.
Proof.
Step 1 (scaled interpolation). Write the recursion in signed-noise form $\theta_{t+1}=\theta_{t}-\eta h(\theta_{t})+\eta\,\zeta_{t+1}$ with $\zeta_{t+1}:=-\xi_{t}$, so $(\zeta_{t+1},\mathcal{F}_{t+1})$ is a martingale difference. Define the piecewise-linear interpolation $\bar{\theta}^{\eta}(s)$ on algorithmic time $s=t\eta$ by $\bar{\theta}^{\eta}(t\eta)=\theta_{t}$ with linear interpolation.
Step 2 (drift convergence). By local Lipschitzness of $h$ (Assumption 5.1) and the localization condition, $\sum_{k<s/\eta}\eta\,h(\theta_{k})\to\int_{0}^{s}h(\theta(u))\,du$ uniformly on compacts, by standard ODE-approximation theory (Kushner and Yin, 2003, Chapter 5).
Step 3 (predictable quadratic variation). Define $M^{\eta}(s):=\sum_{k<s/\eta}\eta\,\zeta_{k+1}$. By the alignment theorem (Theorem 4.3), $\mathbb{E}[\zeta_{k+1}\zeta_{k+1}^{\top}\mid\mathcal{F}_{k}]=\frac{1}{b}\,G^{\star}(\theta_{k})+R_{k}$ with $\|R_{k}\|_{\operatorname{op}}\leq(c/b)\rho_{k}$ and $\rho_{k}\to 0$ locally. Using $\eta^{2}/b=\eta\tau$:
$\langle M^{\eta}\rangle(s)=\sum_{k<s/\eta}\eta^{2}\mathbb{E}[\zeta_{k+1}\zeta_{k+1}^{\top}\mid\mathcal{F}_{k}]=\tau\sum_{k<s/\eta}\eta\,G^{\star}(\theta_{k})+o(1)\;\to\;\tau\!\int_{0}^{s}\!G^{\star}(\theta(u))\,du$
in probability, uniformly on $[0,S]$.
Step 4 (Lindeberg and tightness). The fourth moment bound (Assumption 3.8(b)) yields $\sum_{k<S/\eta}\mathbb{E}[\|\eta\zeta_{k+1}\|^{2}\mathbf{1}\{\|\eta\zeta_{k+1}\|>\varepsilon\}\mid\mathcal{F}_{k}]\leq CS\eta^{3}m_{4}/(\varepsilon^{2}b^{2})\to 0$, verifying the conditional Lindeberg condition. Tightness of $\{\bar{\theta}^{\eta}\}$ in $D([0,S];\mathbb{R}^{d})$ follows from Aldous’ criterion and the Lipschitz-controlled drift.
Step 5 (martingale FCLT). By the martingale central limit theorem (Ethier and Kurtz, 1986, Theorem 7.1.4), $M^{\eta}(\cdot)\Rightarrow M(\cdot)$ weakly in $D([0,S];\mathbb{R}^{d})$ where $M$ is a continuous martingale with quadratic variation $\tau\int_{0}^{s}G^{\star}(\theta(u))\,du$. Lévy’s characterization gives a Brownian motion $W$ such that $M(s)=\int_{0}^{s}\sqrt{\tau}\,C^{\star}(\theta(u))\,dW_{u}$. Combining drift and martingale convergence identifies the limit as the unique solution of (27). ∎
5.3 The OU regime near a nondegenerate critical point
Assume $\theta^{\star}$ is a stationary point: $h(\theta^{\star})=0$, and define the local curvature $H^{\star}:=\nabla h(\theta^{\star})=\nabla^{2}L(\theta^{\star})$. Let $\Delta_{s}:=\theta_{s}-\theta^{\star}$ denote the centered diffusion.
Assumption 5.5 (Nondegeneracy at $\theta^{\star}$).
$H^{\star}\succ 0$ and $G^{\star}(\theta^{\star})\succ 0$.
Theorem 5.6 (OU linearization).
Under Assumption 5.5 and the conditions of Theorem 5.4, the diffusion (27) admits the local expansion
$d\Delta_{s}=-H^{\star}\Delta_{s}\,ds+\sqrt{\tau}\,C^{\star}(\theta^{\star})\,dW_{s}+\rho_{1}(\Delta_{s})\,ds+\sqrt{\tau}\,\rho_{2}(\Delta_{s})\,dW_{s},$ (28)
where $\|\rho_{1}(\Delta)\|=o(\|\Delta\|)$ and $\|\rho_{2}(\Delta)\|_{\operatorname{op}}=o(1)$ as $\Delta\to 0$. Neglecting the higher-order remainder terms yields the OU SDE
$d\Delta_{s}=-H^{\star}\Delta_{s}\,ds+\sqrt{\tau}\,C^{\star}(\theta^{\star})\,dW_{s},\qquad C^{\star}(\theta^{\star}){C^{\star}(\theta^{\star})}^{\top}=G^{\star}(\theta^{\star}).$ (29)
Proof.
A Taylor expansion gives $h(\theta^{\star}+\Delta)=H^{\star}\Delta+r(\Delta)$ with $\|r(\Delta)\|=o(\|\Delta\|)$. Continuity of $G^{\star}$ implies we can choose a continuous square root $C^{\star}(\theta)$ locally (e.g., the symmetric PSD square root), so
$C^{\star}(\theta^{\star}+\Delta)=C^{\star}(\theta^{\star})+R_{C}(\Delta),\qquad\|R_{C}(\Delta)\|_{\operatorname{op}}=o(1)\quad(\Delta\to 0).$
Substituting into (27) and collecting terms yields (28) with $\rho_{1}(\Delta):=-r(\Delta)$ and $\rho_{2}(\Delta):=R_{C}(\Delta)$. ∎
Corollary 5.7 (Stationary covariance and the Lyapunov equation).
If $H^{\star}\succ 0$, the OU process (29) is ergodic with stationary law $\mathcal{N}(0,\Sigma_{\infty})$. Since $H^{\star}\succ 0$, the matrix $H^{\star}$ is Hurwitz, so the continuous Lyapunov equation
$H^{\star}\Sigma_{\infty}+\Sigma_{\infty}(H^{\star})^{\top}=\tau\,G^{\star}(\theta^{\star})$ (30)
admits a unique symmetric positive definite solution $\Sigma_{\infty}$. In particular, the stationary covariance scales linearly with temperature $\tau=\eta/b$.
Because $H^{\star}$ is symmetric positive definite, left-multiplying (30) by $(H^{\star})^{-1}$ and taking traces gives the trace identity
$\mathrm{Tr}(\Sigma_{\infty})+\mathrm{Tr}\!\bigl((H^{\star})^{-1}\Sigma_{\infty}H^{\star}\bigr)=\tau\,\mathrm{Tr}\!\bigl((H^{\star})^{-1}G^{\star}(\theta^{\star})\bigr),$ (31)
which reduces to $2\,\mathrm{Tr}(\Sigma_{\infty})=\tau\,\mathrm{Tr}(G^{\star}H^{\star-1})$ when $H^{\star}$ and $\Sigma_{\infty}$ commute (e.g., when the eigenbases of curvature and noise align).
Proof.
Write $B:=\sqrt{\tau}\,(G^{\star}(\theta^{\star}))^{1/2}$, so $BB^{\top}=\tau G^{\star}(\theta^{\star})$. Since $^{\star}\succ 0, H-H^{\star}$ is Hurwitz and the OU process is exponentially ergodic with stationary law $\mathcal{N}(0,\Sigma_{\infty})$. Apply Itô’s formula to $\Delta_{s}\Delta_{s}^{\top}$: the drift contributes $(H^{\star}\Sigma_{s}+\Sigma_{s}H^{\star\top})$ and the quadratic variation contributes $BB^{\top}$, giving $\frac{d}{ds}\Sigma_{s}=-H^{\star}\Sigma_{s}-\Sigma_{s}H^{\star\top}+BB^{\top}$. Setting $\frac{d}{ds}\Sigma_{s}=0$ yields (30). Uniqueness follows because for Hurwitz $H^{\star}$ the integral $\Sigma_{\infty}=\int_{0}^{\infty}e^{-H^{\star}s}\,(\tau G^{\star})\,e^{-H^{\star\top}s}\,ds$ converges and is the unique solution (Lemma A.3). ∎
Remark 5.8 (Likelihood case (Fisher-aligned diffusion)).
If $\ell(\theta;z)=-\log p_{\theta}(z)$ and the model is correctly specified, then $G^{\star}(\theta^{\star})=F^{\star}(\theta^{\star})$ and the OU diffusion reads
$d\Delta_{s}=-H^{\star}\Delta_{s}\,ds+\sqrt{\tau}\,C_{F}^{\star}\,dW_{s},\qquad C_{F}^{\star}{C_{F}^{\star}}^{\top}=F^{\star}(\theta^{\star}),$
making Fisher geometry explicit in both transient fluctuations and in the stationary Lyapunov balance (30).
5.4 Discrete-time stationary covariance: an exact bridge
The OU viewpoint is not the only consequence of the covariance identification. The same $G^{\star}$ governs the exact linearized discrete-time stationary covariance; the continuous-time Lyapunov equation emerges as the small-stepsize asymptotic reduction. This subsection makes the bridge explicit, so that the paper’s conclusions do not rest on the diffusion approximation alone.
Proposition 5.9 (Discrete-time linearized stationary covariance).
Consider the linearized discrete-time recursion around $\theta^{\star}$ with constant stepsize $\eta$ and batch size $b$:
$\Delta_{t+1}=(I-\eta H^{\star})\,\Delta_{t}+\eta\,\zeta_{t+1},\qquad\mathbb{E}[\zeta_{t+1}\zeta_{t+1}^{\top}\mid\mathcal{F}_{t}]=\frac{1}{b}\,G^{\star}(\theta^{\star})+r_{t},$ (32)
where $\|r_{t}\|_{\operatorname{op}}=o(b^{-1})+O(n^{-1})$ as in (25). If $\rho(I-\eta H^{\star})<1$ (equivalently, $\eta<2/\lambda_{\max}(H^{\star})$), then the iterates are geometrically ergodic and the stationary covariance $\Sigma_{\eta}:=\lim_{t\to\infty}\mathrm{Cov}(\Delta_{t})$ satisfies the discrete Lyapunov equation
$\Sigma_{\eta}=(I-\eta H^{\star})\,\Sigma_{\eta}\,(I-\eta H^{\star})^{\top}+\frac{\eta^{2}}{b}\,G^{\star}(\theta^{\star}),$ (33)
up to the same local remainder terms already used in the diffusion analysis.
Proof.
From (32), compute $\mathrm{Cov}(\Delta_{t+1})=(I-\eta H^{\star})\mathrm{Cov}(\Delta_{t})(I-\eta H^{\star})^{\top}+\eta^{2}\mathbb{E}[\zeta_{t+1}\zeta_{t+1}^{\top}\mid\mathcal{F}_{t}]$, using $\mathbb{E}[\zeta_{t+1}\mid\mathcal{F}_{t}]=0$. Under $\rho(I-\eta H^{\star})<1$, the map $\Sigma\mapsto(I-\eta H^{\star})\Sigma(I-\eta H^{\star})^{\top}+Q$ is a contraction on positive semidefinite matrices, so the iteration converges to the unique fixed point (33). The spectral radius condition $\rho(I-\eta H^{\star})<1$ is equivalent to $0<\eta<2/\lambda_{\max}(H^{\star})$, since $H^{\star}\succ 0$. ∎
Remark 5.10 (Continuous-time reduction and discrete-time corrections).
Expanding (33) and dividing by $\eta$, one recovers $H^{\star}\Sigma_{\eta}+\Sigma_{\eta}H^{\star\top}=\frac{\eta}{b}G^{\star}+\eta H^{\star}\Sigma_{\eta}H^{\star\top}$. Sending $\eta\downarrow 0$ with $\tau=\eta/b$ fixed drops the $O(\eta)$ correction term and yields the continuous Lyapunov equation (30). The correction is governed by $\eta\lambda_{\max}(H^{\star})$: when this product is small the two benchmarks nearly coincide; for moderate stepsizes the discrete equation is the correct comparator (Section 8). The covariance identification $\mathrm{Cov}(\xi_{t})\approx b^{-1}G^{\star}$ is logically prior to both analyses: the discrete Lyapunov equation is its exact linearized consequence, and the OU equation is the small-$\eta$ reduction.
6 Convergence Rates in Fisher Geometry
This section establishes matching upper and lower bounds for SGD when the error is measured in a frozen Fisher metric at a local optimum $\theta^{\star}$. Once the mini-batch noise covariance is identified (Section 4), the remaining work is a localized stochastic-approximation argument in a fixed geometry, followed by an information-theoretic lower bound expressed in the same metric. Convention. All rates in this section are stated in Fisher/Godambe risk $\mathbb{E}[\theta_{T}^{\top}F^{\star}\theta_{T}]$ unless explicitly noted otherwise; any Euclidean norm that appears is immediately converted.
6.1 Local Fisher Geometry and a Frozen-Metric Regime
Work in a neighborhood of $\theta^{\star}$ where the Fisher information varies slowly. For clean rate statements we freeze the metric
$F^{\star}:=F(\theta^{\star}),\qquad\|v\|_{F^{\star}}^{2}:=v^{\top}F^{\star}v.$
(Allowing a slowly varying metric $F(\theta)$ is standard via localization; the frozen-metric bounds below are the local core and are what we use later for oracle complexity.)
Lemma 6.1 (Sufficient condition for the frozen-metric regime).
Suppose $F(\theta)$ is Lipschitz on a neighborhood $\mathcal{U}$ of $\theta^{\star}$: $\|F(\theta)-F(\theta^{\star})\|_{\operatorname{op}}\leq L_{F}\|\theta-\theta^{\star}\|_{2}$ for all $\theta\in\mathcal{U}$. If the SGD iterates satisfy $\|\theta_{t}-\theta^{\star}\|_{2}=o(1)$ almost surely, then for every $\epsilon>0$ there exists $t_{0}$ such that for $t\geq t_{0}$, $(1-\epsilon)F^{\star}\preceq F(\theta_{t})\preceq(1+\epsilon)F^{\star}$.
Let $h(\theta):=\nabla L(\theta)$ and decompose the mini-batch estimator as
$g_{t}(\theta_{t})=h(\theta_{t})+\xi_{t},\qquad\mathbb{E}[\xi_{t}\mid\mathcal{F}_{t}]=0,$
where $\mathcal{F}_{t}$ is the natural filtration. Under Fisher alignment (Section 4), the conditional covariance satisfies the local approximation
$\mathbb{E}[\xi_{t}\xi_{t}^{\top}\mid\mathcal{F}_{t}]\;=\;\frac{1}{b}\,F^{\star}\;+\;r_{t},\qquad\|r_{t}\|_{\operatorname{op}}\leq\frac{c_{r}}{b}\,\rho_{t},$ (34)
with $\rho_{t}\to 0$ in the local regime.
Define the Fisher-strong-convexity constant at $\theta^{\star}$:
$\mu_{F}\;:=\;\lambda_{\min}\!\left(\frac{(F^{\star})^{1/2}H^{\star}(F^{\star})^{-1/2}+(F^{\star})^{-1/2}H^{\star}(F^{\star})^{1/2}}{2}\right),\qquad H^{\star}:=\nabla^{2}L(\theta^{\star}).$ (35)
Assumption 6.2 (Local coercivity in the frozen Fisher geometry).
The constant $\mu_{F}$ defined in (35) is strictly positive. This is not automatic from $H^{\star}\succ 0$ and $F^{\star}\succ 0$ alone; it is an explicit compatibility condition between the loss curvature and noise geometry. It holds whenever $H^{\star}$ and $F^{\star}$ are jointly diagonalizable, and more generally whenever $H^{\star}$ and $F^{\star}$ are sufficiently aligned.
Define the effective dimension
$d_{\operatorname{eff}}(F^{\star})\;:=\;\frac{\mathrm{Tr}(F^{\star})}{\|F^{\star}\|_{\operatorname{op}}}\;=\;\frac{\mathrm{Tr}(F^{\star})}{\lambda_{\max}(F^{\star})}\;\in(0,d].$ (36)
Remark 6.3 (Effective dimension as stable rank).
The quantity $d_{\operatorname{eff}}(F^{\star})$ is exactly the stable rank of $F^{\star}$ (Tropp, 2015): $d_{\operatorname{eff}}\leq d$ always, with equality iff $F^{\star}$ is isotropic, and $d_{\operatorname{eff}}\to 1$ when a single direction dominates. It replaces $d$ in both the upper bound (Theorem 6.5) and the oracle complexity (Theorem 7.3), so that when Fisher information is concentrated ($d_{\operatorname{eff}}\ll d$), the guarantees are correspondingly tighter than classical Euclidean bounds.
6.2 Upper Bound
Lemma 6.4 (One-step Fisher Lyapunov drift).
Let $\Delta_{t}:=\theta_{t}-\theta^{\star}$ and $V_{t}:=\|\Delta_{t}\|_{F^{\star}}^{2}$. Assume: (i) $L$ is locally $M$-smooth near $\theta^{\star}$; (ii) $H^{\star}\succ 0$; (iii) (34) holds. Then, as long as $\theta_{t}$ stays in the local neighborhood,
$\mathbb{E}[V_{t+1}\mid\mathcal{F}_{t}]\;\leq\;\Bigl(1-2\mu_{F}\eta_{t}+c_{0}\eta_{t}^{2}\Bigr)V_{t}\;+\;\eta_{t}^{2}\,\frac{1}{b}\,\mathrm{Tr}\!\bigl((F^{\star})^{2}\bigr)\;+\;\eta_{t}^{2}\,\frac{c_{1}}{b}\,\mathrm{Tr}(F^{\star})\,\rho_{t},$ (37)
where $c_{0},c_{1}$ depend only on $(M,\|F^{\star}\|_{\operatorname{op}},c_{r})$.
Proof.
From $\theta_{t+1}=\theta_{t}-\eta_{t}(h(\theta_{t})+\xi_{t})$, $\Delta_{t+1}=\Delta_{t}-\eta_{t}h(\theta_{t})-\eta_{t}\xi_{t}$. Expanding $V_{t+1}=\Delta_{t+1}^{\top}F^{\star}\Delta_{t+1}$ and taking conditional expectation (using $\mathbb{E}[\xi_{t}\mid\mathcal{F}_{t}]=0$) yields
$\mathbb{E}[V_{t+1}\mid\mathcal{F}_{t}]=V_{t}-2\eta_{t}\Delta_{t}^{\top}F^{\star}h(\theta_{t})+\eta_{t}^{2}\,h(\theta_{t})^{\top}F^{\star}h(\theta_{t})+\eta_{t}^{2}\,\mathrm{Tr}\!\Bigl(F^{\star}\,\mathbb{E}[\xi_{t}\xi_{t}^{\top}\mid\mathcal{F}_{t}]\Bigr).$
Using $h(\theta^{\star}+\Delta)=H^{\star}\Delta+o(\|\Delta\|)$, we have $\Delta_{t}^{\top}F^{\star}h(\theta_{t})=\Delta_{t}^{\top}F^{\star}H^{\star}\Delta_{t}+o(V_{t})$. Writing $x_{t}:=(F^{\star})^{1/2}\Delta_{t}$ gives $\Delta_{t}^{\top}F^{\star}H^{\star}\Delta_{t}=x_{t}^{\top}\,\operatorname{Sym}\bigl((F^{\star})^{1/2}H^{\star}(F^{\star})^{-1/2}\bigr)\,x_{t}\geq\mu_{F}V_{t}$. Local $M$-smoothness implies $h(\theta_{t})^{\top}F^{\star}h(\theta_{t})\leq c_{0}V_{t}$. For the noise term, plug (34):
$\mathrm{Tr}\!\Bigl(F^{\star}\,\mathbb{E}[\xi_{t}\xi_{t}^{\top}\mid\mathcal{F}_{t}]\Bigr)=\frac{1}{b}\mathrm{Tr}((F^{\star})^{2})+\mathrm{Tr}(F^{\star}r_{t}).$
Finally $|\mathrm{Tr}(F^{\star}r_{t})|\lesssim\mathrm{Tr}(F^{\star})\|r_{t}\|_{\operatorname{op}}\leq(c_{1}/b)\mathrm{Tr}(F^{\star})\rho_{t}$, giving (37). ∎
Theorem 6.5 (Upper bound in frozen Fisher geometry).
Assume Lemma 6.4. Take $\eta_{t}=\eta_{0}/t$ with $\eta_{0}>1/(2\mu_{F})$, and assume $\sum_{t\geq 1}\eta_{t}^{2}\rho_{t}<\infty$. Then for all $T\geq 2$,
$\mathbb{E}\!\left[\|\theta_{T}-\theta^{\star}\|_{F^{\star}}^{2}\right]\;\leq\;\frac{C_{1}}{b}\cdot\frac{\mathrm{Tr}\!\bigl((F^{\star})^{2}\bigr)}{\mu_{F}}\cdot\frac{1}{T}\;+\;C_{2}\cdot\frac{\|\theta_{1}-\theta^{\star}\|_{F^{\star}}^{2}}{T^{2\mu_{F}\eta_{0}}}\;+\;\frac{C_{3}}{b}\cdot\frac{\mathrm{Tr}(F^{\star})}{T},$ (38)
where $C_{1},C_{2},C_{3}$ depend only on $(\eta_{0},\mu_{F},c_{0},c_{1})$.
Moreover, using $\mathrm{Tr}((F^{\star})^{2})\leq\lambda_{\max}(F^{\star})\,\mathrm{Tr}(F^{\star})$, we obtain
$\mathbb{E}\!\left[\|\theta_{T}-\theta^{\star}\|_{F^{\star}}^{2}\right]$ $\;\leq\;\frac{\widetilde{C}_{1}}{b}\cdot\frac{\lambda_{\max}(F^{\star})\,\mathrm{Tr}(F^{\star})}{\mu_{F}}\cdot\frac{1}{T}\;+\;\widetilde{C}_{2}\cdot\frac{\|\theta_{1}-\theta^{\star}\|_{F^{\star}}^{2}}{T^{2\mu_{F}\eta_{0}}}$
$\;=\;\frac{\widetilde{C}_{1}}{b}\cdot\frac{\lambda_{\max}(F^{\star})^{2}}{\mu_{F}}\cdot\frac{d_{\operatorname{eff}}(F^{\star})}{T}\;+\;\widetilde{C}_{2}\cdot\frac{\|\theta_{1}-\theta^{\star}\|_{F^{\star}}^{2}}{T^{2\mu_{F}\eta_{0}}}.$ (39)
Proof.
Take total expectations in (37) to obtain $\mathbb{E}[V_{t+1}]\leq a_{t}\,\mathbb{E}[V_{t}]+u_{t}$ with $a_{t}:=1-2\mu_{F}\eta_{t}+c_{0}\eta_{t}^{2}$ and $u_{t}:=\eta_{t}^{2}\frac{1}{b}\mathrm{Tr}((F^{\star})^{2})+\eta_{t}^{2}\frac{c_{1}}{b}\mathrm{Tr}(F^{\star})\rho_{t}$. For $\eta_{t}=\eta_{0}/t$, there exists $t_{0}$ such that for $t\geq t_{0}$, $a_{t}\leq 1-\mu_{F}\eta_{0}/t$. Define $A_{t,T}:=\prod_{s=t}^{T-1}a_{s}$. Unrolling gives $\mathbb{E}[V_{T}]\leq A_{1,T}V_{1}+\sum_{t=1}^{T-1}A_{t+1,T}u_{t}$. Standard product bounds yield $A_{1,T}\leq CT^{-2\mu_{F}\eta_{0}}$ and $A_{t+1,T}\leq C(t/T)^{\mu_{F}\eta_{0}}$. Then, since $\mu_{F}\eta_{0}>1/2$, $\sum_{t\leq T}A_{t+1,T}\eta_{t}^{2}=O(1/T)$, giving the leading $(\mathrm{Tr}((F^{\star})^{2})/b)(1/T)$ term. The remainder term is controlled because $\sum_{t}\eta_{t}^{2}\rho_{t}<\infty$ and $A_{t+1,T}\leq 1$, yielding an additional $O((\mathrm{Tr}(F^{\star})/b)(1/T))$ contribution. This proves (38). The reduction to (6.5) follows from $\mathrm{Tr}((F^{\star})^{2})\leq\lambda_{\max}(F^{\star})\mathrm{Tr}(F^{\star})$. ∎
Remark 6.6 (Consistency with the OU viewpoint).
The $1/(Tb)$ scaling in the frozen Fisher norm is consistent with the OU/Lyapunov stationary analysis: both predict the same leading-order dependence on temperature $\tau=\eta/b$ and intrinsic curvature $\mu_{F}$.
6.3 Lower Bound
The lower bound proceeds in three layers of increasing generality. We first state the cleanest result in the parametric likelihood/Fisher case, then interpret it through oracle information, and finally discuss the extension to general losses.
6.3.1 Layer 1: Parametric Fisher-case lower bound.
Assumption 6.7 (Local parametric regularity for van Trees).
Let $\{P_{\theta}:\theta\in\Theta_{\operatorname{loc}}\}$ be a parametric family on a neighborhood $\Theta_{\operatorname{loc}}$ of $\theta^{\star}$ with densities $p_{\theta}$. Assume:
  1. (i)
    $\theta\mapsto p_{\theta}(z)$ is twice continuously differentiable for a.e. $z$;
  2. (ii)
    the Fisher information $F(\theta):=\mathbb{E}_{\theta}[s_{\theta}s_{\theta}^{\top}]$ is continuous and positive definite on $\Theta_{\operatorname{loc}}$;
  3. (iii)
    $\int\|s_{\theta}(z)\|^{2}\,p_{\theta}(z)\,dz$ is uniformly bounded on $\Theta_{\operatorname{loc}}$;
  4. (iv)
    there exists a smooth prior density $\pi$ supported on $\Theta_{\operatorname{loc}}$ with finite Fisher information $J_{\pi}:=\int\|\nabla\log\pi(\theta)\|^{2}\pi(\theta)\,d\theta<\infty$.
Theorem 6.8 (Fisher-case lower bound via van Trees).
Suppose Assumption 6.7 holds and the estimator $\hat{\theta}_{N}$ is based on $N$ i.i.d. observations from $P_{\theta^{\star}}$. Then
$\inf_{\hat{\theta}_{N}}\;\sup_{\theta\in\Theta_{\operatorname{loc}}}\mathbb{E}_{\theta}\!\left[\|\hat{\theta}_{N}-\theta\|_{F(\theta)}^{2}\right]\;\geq\;\frac{c\,d}{N},$ (40)
for a universal constant $c>0$.
Proof.
This is the special case of Proposition 6.10 below with i.i.d. observations (so the oracle information bound holds with equality). We give the self-contained argument in the proof of Proposition 6.10. ∎
6.3.2 Layer 2: Oracle-information interpretation.
The heart of the lower bound is not i.i.d. sampling per se, but bounded predictable information growth. We now restate the result in terms of an abstract oracle transcript.
Assumption 6.9 (Oracle–information structure).
The oracle produces observations $Y_{t}=\psi(\theta_{t};Z_{t})$, where:
  1. (i)
    $\theta_{t}$ is $\mathcal{F}_{t-1}$-measurable, with filtration $\mathcal{F}_{t}:=\sigma(Y_{1},\dots,Y_{t})$;
  2. (ii)
    Conditional unbiasedness: $\mathbb{E}_{\theta^{\star}}[Y_{t}\mid\mathcal{F}_{t-1}]=0$;
  3. (iii)
    Predictable information bound: $\mathbb{E}_{\theta^{\star}}[Y_{t}Y_{t}^{\top}\mid\mathcal{F}_{t-1}]\preceq\mathcal{I}_{1}(\theta^{\star})$, a.s., where $\mathcal{I}_{1}(\theta^{\star}):=F(\theta^{\star})=F^{\star}$ in the likelihood case;
  4. (iv)
    Uniform moment control: $\sup_{t}\mathbb{E}_{\theta^{\star}}\|Y_{t}\|^{4}<\infty$.
Under Assumption 6.9, the predictable quadratic variation of the oracle transcript satisfies $\sum_{t=1}^{N}\mathbb{E}[Y_{t}Y_{t}^{\top}\mid\mathcal{F}_{t-1}]\preceq N\mathcal{I}_{1}(\theta^{\star})$, so the total information available to any $\mathcal{F}_{N}$-measurable estimator is bounded by $N\mathcal{I}_{1}$.
Proposition 6.10 (Lower bound under predictable information budget).
Suppose Assumption 6.7 holds for the parametric family $\{P_{\theta}\}$ and the oracle satisfies Assumption 6.9 with $\mathcal{I}_{1}(\theta^{\star})=F^{\star}$. Additionally, assume that each conditional law of $Z_{t}$ given $\mathcal{F}_{t-1}$ has a density differentiable in $\theta$, so that the conditional Fisher information $I_{t}(\theta\mid\mathcal{F}_{t-1})$ exists and satisfies $I_{t}(\theta^{\star}\mid\mathcal{F}_{t-1})\preceq F^{\star}$ a.s. Then every $\mathcal{F}_{N}$-measurable estimator $\hat{\theta}_{N}$ satisfies
$\inf_{\hat{\theta}_{N}}\;\sup_{\theta\in\Theta_{\operatorname{loc}}}\mathbb{E}_{\theta}\!\left[\|\hat{\theta}_{N}-\theta\|_{F(\theta)}^{2}\right]\;\geq\;\frac{c\,d}{N},$ (41)
for a universal constant $c>0$.
In particular, if one SGD iteration uses a mini-batch of size $b$ (so $N=Tb$ oracle calls), then $\inf_{\hat{\theta}_{T}}\sup_{\theta}\mathbb{E}[\|\hat{\theta}_{T}-\theta\|_{F(\theta)}^{2}]\geq c\,d/(Tb)$.
Proof.
The argument extends Theorem 6.8 from i.i.d. observations to the adaptive oracle transcript. The van Trees bound requires an upper bound on the Fisher information $I_{N}(\theta)$ of the joint likelihood of $(Y_{1},\ldots,Y_{N})$ with respect to $\theta$. Three ingredients deliver this bound.
Step 1 (Adaptivity does not inflate per-step information). At step $t$, the query $\theta_{t}$ is $\mathcal{F}_{t-1}$-measurable, but $Z_{t}\sim P_{\theta}$ is drawn independently of $\mathcal{F}_{t-1}$ given $\theta$. Hence the conditional distribution of $Z_{t}$ given $\mathcal{F}_{t-1}$ depends on $\theta$ only through $\{P_{\theta}\}$, not through the past transcript.
Step 2 (Second-moment bound == per-step Fisher bound). Under correct specification, $Y_{t}=-s_{\theta}(Z_{t})$ is the negative score. The conditional Fisher information of $Z_{t}$ given $\mathcal{F}_{t-1}$ about $\theta$ is $I_{t}(\theta\mid\mathcal{F}_{t-1})=\mathbb{E}_{\theta}[s_{\theta}s_{\theta}^{\top}\mid\mathcal{F}_{t-1}]=F(\theta)$—exactly the score second-moment matrix. So Assumption 6.9(iii) directly bounds $I_{t}(\theta^{\star}\mid\mathcal{F}_{t-1})\preceq F^{\star}$. To extend from $\theta^{\star}$ to all $\theta\in\Theta_{\operatorname{loc}}$, continuity of $F(\theta)$ (Assumption 6.7(ii)) gives $F(\theta)\preceq(1+\epsilon)F^{\star}$ on a sufficiently small $\Theta_{\operatorname{loc}}$; the factor $(1+\epsilon)$ is absorbed into the universal constant $c$.
Step 3 (Chain rule and conclusion). The chain rule for conditional Fisher information applied to the sequential factorization of the joint likelihood gives
$I_{N}(\theta)=\sum_{t=1}^{N}\mathbb{E}_{\theta}\!\bigl[I_{t}(\theta\mid\mathcal{F}_{t-1})\bigr]\preceq N\,(1{+}\epsilon)\,F^{\star}.$
Substituting into van Trees with prior $\pi$ and sending $\epsilon\to 0$ yields (41). ∎
Remark 6.11 (Beyond i.i.d. oracles).
The classical i.i.d. setting satisfies Assumption 6.9 with equality in (iii). The assumption also accommodates exchangeable or conditionally i.i.d. sampling schemes (e.g., de Finetti mixtures) provided the per-step conditional information is bounded by $\mathcal{I}_{1}$. What is required is a predictable information bound on the oracle transcript, not independence across calls. The critical logical step is that the van Trees inequality applies to the joint transcript likelihood, whose total Fisher information is controlled by summing the per-step conditional bounds.
Remark 6.12 (What enters the lower bound).
The lower bound (41) involves $d$ (ambient dimension) and $N$ (oracle budget). The quantities $d_{\mathrm{eff}}$ and $\kappa_{F}$ enter the upper bound and oracle complexity (Theorems 6.5 and 7.3) but not the lower bound, which is stated in terms of $d$ to give the sharpest minimax comparison. When $d_{\mathrm{eff}}\ll d$, the gap between upper and lower bounds tightens correspondingly.
6.3.3 Layer 3: General-loss / Godambe extension.
Remark 6.13 (Extension to general losses).
In the general-loss case, the same lower-bound form $\Omega(d/N)$ is recovered under local asymptotic conditions that identify the relevant sandwich geometry. Specifically, if the loss satisfies a local asymptotic normality (LAN) condition at $\theta^{\star}$ with information matrix $J_{\mathrm{God}}(\theta^{\star})=H^{\star\top}G^{\star-1}H^{\star}$, then the van Trees argument carries through with $F^{\star}$ replaced by $G^{\star}(\theta^{\star})$ and the dual norm adjusted accordingly. The LAN condition is standard for smooth M-estimators under regularity conditions comparable to those already imposed in Section 3; see Le Cam (1986) and van der Vaart (1998). We state this extension as a remark rather than a theorem because the full verification of LAN for the adaptive oracle transcript of Assumption 6.9 requires additional technical conditions (uniform local asymptotic equicontinuity) that we do not develop here.
Corollary 6.14 (Matching rates (up to constants)).
Under Theorem 6.5 and the oracle model of Assumption 6.9,
$\mathbb{E}\!\left[\|\theta_{T}-\theta^{\star}\|_{F^{\star}}^{2}\right]=\Theta\!\left(\frac{1}{Tb}\right)\quad\text{equivalently}\quad\Theta\!\left(\frac{1}{N}\right)\text{ per oracle call},\quad N=Tb.$
7 Oracle Complexity
This section turns the Fisher-geometry mean-square bounds of Section 6 into an explicit oracle-complexity guarantee for achieving $\varepsilon$-stationarity. The steps are standard—smoothness descent, averaged-gradient bound, martingale concentration (cf. Nesterov, 2004, Chapter 2)—but we express the result in the natural Fisher dual norm and keep the statistical conditioning explicit.
Throughout, we work locally near $\theta^{\star}$ and freeze the metric $F^{\star}:=F(\theta^{\star})\succ 0$ (for general losses, replace $F^{\star}$ by $G^{\star}(\theta^{\star})$; see Remark 7.5).
7.1 Stationarity in the Fisher dual norm
Definition 7.1 ($\varepsilon$-stationarity in Fisher dual norm).
Let $F^{\star}\succ 0$. We say $\theta$ is $\varepsilon$-stationary (in the Fisher metric) if
$\|\nabla L(\theta)\|_{(F^{\star})^{-1}}\;:=\;\sqrt{\nabla L(\theta)^{\top}(F^{\star})^{-1}\nabla L(\theta)}\;\leq\;\varepsilon.$
7.2 From Fisher distance to Fisher-dual stationarity
Lemma 7.2 (Distance-to-stationarity conversion).
Assume $L$ is $M$-smooth on a neighborhood $\mathcal{U}$ of $\theta^{\star}$ and $F^{\star}\succ 0$. Then for all $\theta\in\mathcal{U}$,
$\|\nabla L(\theta)\|_{(F^{\star})^{-1}}^{2}\;\leq\;\frac{M^{2}}{\lambda_{\min}(F^{\star})^{2}}\,\|\theta-\theta^{\star}\|_{F^{\star}}^{2}.$ (42)
Moreover, if $\theta\to\theta^{\star}$ and $\nabla L(\theta)=H^{\star}(\theta-\theta^{\star})+o(\|\theta-\theta^{\star}\|)$, then locally
$\|\nabla L(\theta)\|_{(F^{\star})^{-1}}^{2}\;\leq\;(1+o(1))\,\lambda_{\max}\!\bigl(F^{\star-1/2}H^{\star}F^{\star-1/2}\bigr)^{2}\,\|\theta-\theta^{\star}\|_{F^{\star}}^{2}.$ (43)
Proof.
By $M$-smoothness and $\nabla L(\theta^{\star})=0$, $\|\nabla L(\theta)\|_{2}\leq M\|\theta-\theta^{\star}\|_{2}$. The dual-norm bound $\|v\|_{(F^{\star})^{-1}}^{2}\leq\lambda_{\min}(F^{\star})^{-1}\|v\|_{2}^{2}$ converts Euclidean gradient norm to Fisher dual norm, and $\|\theta-\theta^{\star}\|_{2}^{2}\leq\lambda_{\min}(F^{\star})^{-1}\|\theta-\theta^{\star}\|_{F^{\star}}^{2}$ converts Fisher distance to Euclidean. Combining: $\|\nabla L\|_{(F^{\star})^{-1}}^{2}\leq\lambda_{\min}(F^{\star})^{-1}M^{2}\lambda_{\min}(F^{\star})^{-1}\|\theta-\theta^{\star}\|_{F^{\star}}^{2}=M^{2}/\lambda_{\min}(F^{\star})^{2}\cdot\|\theta-\theta^{\star}\|_{F^{\star}}^{2}$. The local refinement follows by replacing $M$ by $\|H^{\star}\|_{\operatorname{op}}$ up to $1+o(1)$ via the linearization. ∎
7.3 High-probability oracle complexity
Theorem 7.3 (Oracle complexity for Fisher-dual stationarity).
Assume the local regularity conditions of Section 3 hold on a neighborhood $\mathcal{U}$ of $\theta^{\star}$, and suppose that on $\mathcal{U}$ the mini-batch noise is Fisher-aligned in the sense that, for some batch size $b\geq 1$,
$\mathbb{E}[\xi_{t}\mid\mathcal{F}_{t}]=0,\qquad\mathbb{E}[\xi_{t}\xi_{t}^{\top}\mid\mathcal{F}_{t}]\preceq\frac{1}{b}\,F^{\star}.$ (44)
Additionally, assume conditional sub-Gaussianity of the noise: for all $v\in\mathbb{R}^{d}$ with $\|v\|=1$ and all $\lambda\in\mathbb{R}$, $\mathbb{E}[\exp(\lambda\,v^{\top}\xi_{t})\mid\mathcal{F}_{t}]\leq\exp(\lambda^{2}\|F^{\star}\|_{\operatorname{op}}/(2b))$ a.s. (This is satisfied, for example, when per-sample gradients are almost surely bounded on $\mathcal{U}$.) Run SGD with constant batch size $b$ and constant stepsize $\eta\leq 1/(4M)$:
$\theta_{t+1}=\theta_{t}-\eta\bigl(\nabla L(\theta_{t})+\xi_{t}\bigr).$
Define the Fisher condition number and intrinsic dimension
$\kappa_{F}:=\frac{\lambda_{\max}(F^{\star})}{\lambda_{\min}(F^{\star})},\qquad d_{\mathrm{eff}}(F^{\star}):=\frac{\mathrm{Tr}(F^{\star})}{\lambda_{\max}(F^{\star})}.$
Then there exists a universal constant $C>0$ such that, for any $\delta\in(0,1)$, with probability at least $1-\delta$,
$\min_{0\leq t\leq T-1}\;\|\nabla L(\theta_{t})\|_{(F^{\star})^{-1}}^{2}\;\leq\;C\left(\frac{L(\theta_{0})-L(\theta^{\star})}{\eta\,T\,\lambda_{\min}(F^{\star})}\;+\;\frac{\eta}{b}\,\kappa_{F}\,d_{\mathrm{eff}}(F^{\star})\right)\log\!\frac{1}{\delta}.$ (45)
In particular, choosing $(\eta,T)$ so that the right-hand side is $\leq\varepsilon^{2}$ implies $\min_{t<T}\|\nabla L(\theta_{t})\|_{(F^{\star})^{-1}}\leq\varepsilon$ with probability at least $1-\delta$.
Equivalently, in terms of oracle calls $N:=Tb$ (sample-gradient evaluations), one sufficient scaling (up to constants and ignoring the transient term $\frac{L(\theta_{0})-L(\theta^{\star})}{\eta T\lambda_{\min}(F^{\star})}$) is
$N\;=\;\Theta\!\left(\frac{\kappa_{F}\,d_{\mathrm{eff}}(F^{\star})}{\varepsilon^{2}}\,\log\!\frac{1}{\delta}\right).$ (46)
Proof.
Step 1 (descent inequality). By $M$-smoothness and $\eta\leq 1/(4M)$, expand $L(\theta_{t+1})\leq L(\theta_{t})+\langle\nabla L,\theta_{t+1}-\theta_{t}\rangle+\frac{M}{2}\|\theta_{t+1}-\theta_{t}\|^{2}$ with $\theta_{t+1}-\theta_{t}=-\eta(\nabla L(\theta_{t})+\xi_{t})$. After expanding and using $\eta\leq 1/(4M)$, so $1-M\eta/2\geq 1-1/8=7/8$ and $\eta M\leq 1/4$:
$L(\theta_{t+1})\leq L(\theta_{t})-\tfrac{3\eta}{4}\|\nabla L(\theta_{t})\|_{2}^{2}+\eta(1+M\eta)\langle\nabla L(\theta_{t}),\xi_{t}\rangle+\tfrac{M\eta^{2}}{2}\|\xi_{t}\|_{2}^{2}.$
Step 2 (define martingale). Let $\Delta M_{t+1}:=\eta\langle\nabla L(\theta_{t}),\xi_{t}\rangle$ and $M_{T}:=\sum_{t=0}^{T-1}\Delta M_{t+1}$; since $\mathbb{E}[\xi_{t}\mid\mathcal{F}_{t}]=0$, this is a martingale.
Step 3 (sum the descent). Summing from $t=0$ to T1T-1:
$\tfrac{3\eta}{4}\,S_{T}\;\leq\;\Delta L+C_{1}M_{T}+\tfrac{M\eta^{2}}{2}\,Q_{T},\quad\Delta L:=L(\theta_{0})-L(\theta^{\star}),$
where $Q_{T}:=\sum_{t<T}\|\xi_{t}\|_{2}^{2}$. To replace the random $Q_{T}$ by a deterministic bound we need concentration, not just conditional moments. Under conditional sub-Gaussianity, $\|\xi_{t}\|_{2}^{2}$ is conditionally sub-exponential with $\mathbb{E}[\|\xi_{t}\|_{2}^{2}\mid\mathcal{F}_{t}]\leq b^{-1}\mathrm{Tr}(F^{\star})$; Bernstein’s inequality for sums of conditionally sub-exponential random variables (see, e.g., Wainwright 2019, Cor. 2.21) gives, with probability at least $1-\delta$, $Q_{T}\leq(T/b)\mathrm{Tr}(F^{\star})+c\sqrt{T\,d}\,(\|F^{\star}\|_{\operatorname{op}}/b)\sqrt{\log(1/\delta)}.$ Since $M\eta^{2}/2\leq\eta/8$ (as $\eta\leq 1/(4M)$), the deviation is $O(\eta\sqrt{T/b}\,\sqrt{d}\,\|F^{\star}\|_{\operatorname{op}}\sqrt{\log(1/\delta)}\,/b)$, which is lower-order relative to the bias term $\Delta L/(\eta T)$ in the final bound. Absorbing it into the constant $C$ (and taking a union bound over the two $\delta$-events, i.e. replacing $\delta$ by $\delta/2$ throughout):
$\tfrac{3\eta}{4}\,S_{T}\;\leq\;\Delta L+C_{1}M_{T}+\tfrac{M\eta^{2}T}{2b}\,\mathrm{Tr}(F^{\star})\log\!\tfrac{2}{\delta},$
with probability at least $1-\delta$.
Step 4 (sub-Gaussian martingale concentration). Each increment satisfies $\Delta M_{t+1}=\eta\langle\nabla L(\theta_{t}),\xi_{t}\rangle$ where $\nabla L(\theta_{t})$ is $\mathcal{F}_{t}$-measurable. By the conditional sub-Gaussianity assumption, the scalar $\langle v,\xi_{t}\rangle$ with $v:=\nabla L(\theta_{t})/\|\nabla L(\theta_{t})\|_{2}$ has sub-Gaussian parameter $\sigma_{t}^{2}\leq\|F^{\star}\|_{\operatorname{op}}/b$, so $\Delta M_{t+1}$ is conditionally sub-Gaussian with parameter $c_{t}^{2}:=\eta^{2}\|\nabla L(\theta_{t})\|_{2}^{2}\|F^{\star}\|_{\operatorname{op}}/b$. By the standard exponential supermartingale bound for sums of conditionally sub-Gaussian increments (see, e.g., Wainwright 2019, Theorem 2.19), with probability at least $1-\delta$, $M_{T}\leq\sqrt{2\sum_{t<T}c_{t}^{2}\,\log(1/\delta)}=\eta\sqrt{2(\|F^{\star}\|_{\operatorname{op}}/b)\,S_{T}\,\log(1/\delta)}$, where $S_{T}:=\sum_{t<T}\|\nabla L(\theta_{t})\|_{2}^{2}$.
Step 5 (solve for gradient bound). Substituting the Step 4 tail bound into Step 3 gives a quadratic inequality in $\sqrt{S_{T}}$; solving:
$\min_{t<T}\|\nabla L(\theta_{t})\|_{2}^{2}\;\leq\;\frac{S_{T}}{T}\;\leq\;C\!\left(\frac{\Delta L}{\eta T}+\frac{\eta}{b}\mathrm{Tr}(F^{\star})\right)\!\log\!\frac{1}{\delta}.$
Step 6 (Fisher dual norm). Use $\|v\|_{(F^{\star})^{-1}}^{2}\leq\lambda_{\min}(F^{\star})^{-1}\|v\|_{2}^{2}$ and $\mathrm{Tr}(F^{\star})/\lambda_{\min}(F^{\star})=\kappa_{F}\,d_{\operatorname{eff}}(F^{\star})$:
$\min_{t<T}\|\nabla L(\theta_{t})\|_{(F^{\star})^{-1}}^{2}\;\leq\;C\!\left(\frac{\Delta L}{\eta T\lambda_{\min}(F^{\star})}+\frac{\eta}{b}\,\kappa_{F}\,d_{\operatorname{eff}}(F^{\star})\right)\!\log\!\frac{1}{\delta},$
establishing (45). The oracle complexity (46) follows by balancing the two terms and setting $N=Tb$. ∎
Remark 7.4 (Statistical vs. Euclidean conditioning).
Classical oracle-complexity results use the Hessian condition number $\kappa_{H}$ and ambient dimension $d$; here they are replaced by the Fisher counterparts $\kappa_{F}$ and $d_{\operatorname{eff}}$. The substitution is not cosmetic: a problem may be Euclidean-stiff ($\kappa_{H}\gg 1$) yet statistically well-conditioned ($\kappa_{F}=O(1)$) when curvature and information share the same eigenstructure; conversely, isotropic curvature does not guarantee favorable statistical complexity.
Remark 7.5 (General losses / Godambe geometry).
For general losses, replace $F^{\star}$ by $G^{\star}(\theta^{\star})$ throughout and interpret the dual norm accordingly. The alignment condition (44) is then exactly the local conclusion of Theorem 4.3.
Remark 7.6 (Iteration vs. oracle complexity).
Equation (46) counts oracle calls $N=Tb$. If reporting iteration complexity, keep $b$ explicit and interpret $T=N/b$.
Remark 7.7 (Interpreting the complexity constants).
The product $\kappa_{F}\,d_{\operatorname{eff}}$ in (46) measures the total cost of resolving all statistically relevant directions at the hardest direction’s scale (see Remark 6.3 for the role of $d_{\operatorname{eff}}$ as the stable rank of $F^{\star}$). Equation (46) links the sampling budget NN to achievable stationarity through intrinsic problem geometry.
Corollary 7.8 (Batch-size scaling under a fixed sample budget).
Fix a total sampling budget $N$ and run $T=N/b$ iterations of constant-stepsize SGD with batch size $b$. Equating the bias term $\sim(L(\theta_{0})-L^{\star})/(\eta T\lambda_{\min}(F^{\star}))$ and the variance term $\sim\eta\kappa_{F}d_{\operatorname{eff}}/b$ in (45) yields the balanced stepsize $\eta^{\star}\propto\sqrt{b/(T\kappa_{F}d_{\operatorname{eff}})}$. Substituting $T=N/b$, the resulting Fisher-dual stationarity gap satisfies
$\min_{t<T}\|\nabla L(\theta_{t})\|_{(F^{\star})^{-1}}^{2}\;=\;\mathcal{O}\!\left(\sqrt{\frac{\kappa_{F}\,d_{\operatorname{eff}}}{N}}\right),$
which is independent of $b$ at leading order. The practical implication for OR settings is that, under this budget-constrained regime, small batches lose nothing in worst-case guarantee while gaining more frequent iterates; the Fisher/Godambe geometry determines the constant but not the scaling in $N$.
8 Numerical Validation
We validate the theory through six diagnostic experiments summarized in Table 3. The experiments probe downstream consequences of the identified covariance $\mathrm{Cov}(g_{B})\approx b^{-1}G^{\star}$ (Theorem 4.3): if the identification is correct, the Lyapunov equilibrium, convergence rates, and directional variance structure must all agree with the predictions derived from $G^{\star}$. Appendix B additionally validates the covariance identity itself under misspecification and approximate exchangeability.
Table 3: Reproducibility capsule: key experimental parameters.
Experiment $d$ Replicates $\eta$ Steps Sweep variable
1 (Batch scaling) 2 500 0.05 200 000 $b\in\{1,2,4,\ldots,512\}$
2 (Angle sweep) 2 500 0.05 200 000 $\theta\in[0,\pi/2]$
3 (Rotation diagnostic) 2 $\phi\in[0,\pi/2]$ (theory)
4 (OU plateau, d=10d\!=\!10) 10 2 000 $0.5/\lambda_{\max}$ 300–10 000 $T$ (horizon)
5 ($1/N$ rate) 10 800 $\eta_{0}/(t\!+\!t_{0})$ 600–9 600 $N=Tb$
6 (Directional var.) 10 800 $\eta_{0}/(t\!+\!t_{0})$ 600–9 600 eigenvectors of $G^{\star}$
Experiments 1–3: Euler–Maruyama, $d=2$. Experiments 4–6: direct SGD recursion, $d=10$, $b=64$.
8.1 Claim 1: The Lyapunov Equilibrium Law Is Quantitatively Predictive
The first claim is that the Fisher–Lyapunov balance (Corollary 5.7) is not merely a qualitative insight but a quantitatively accurate prediction of steady-state risk, and that the shape of the noise covariance—not just its scale—governs which directions receive noise energy at equilibrium. Three experiments validate this at increasing dimensionality.
Throughout, when the simulator uses finite-step (Euler–Maruyama or direct SGD) dynamics, the mathematically correct benchmark is the discrete Lyapunov equation (Proposition 5.9); the continuous-time OU prediction is its small-$\eta$ asymptotic reduction. The regime of approximation accuracy is governed by the product $\eta\lambda_{\max}(H^{\star})$: when this quantity is small (say $\leq 0.1$), the continuous and discrete benchmarks are nearly indistinguishable; for moderate stepsizes the discrete equation is the appropriate comparator.
8.1.1 Experiment 1: Fisher–Lyapunov Equilibrium and $1/b$ Scaling (OU Benchmark).
We begin with the local OU surrogate implied by the diffusion limit near a nondegenerate optimum:
$d\theta_{t}\;=\;-\eta\,H\,\theta_{t}\,dt\;+\;\sqrt{\frac{2\eta}{b}}\;F^{1/2}\,dW_{t},$ (47)
with a two-dimensional sharp/flat curvature $H=\mathrm{diag}(1,0.1)$ and isotropic noise $F=I_{2}$, sweeping $b$ over a wide range. At stationarity, the covariance $\Sigma(b)$ solves the Lyapunov equation
$H\Sigma+\Sigma H\;=\;\frac{2}{b}\,F,$ (48)
so every stationary marginal variance scales as $1/b$; $\eta$ governs mixing time but not the fixed point. For the Euler–Maruyama simulation used here, the discrete Lyapunov equation (49) is the exact benchmark; the continuous equation (48) coincides at leading order because the stepsize satisfies $\eta\lambda_{\max}(H)\leq 0.05$.
Figure 2 overlays simulated stationary variances with the Lyapunov prediction and shows clean $1/b$ scaling across the batch sweep. The flat coordinate mixes more slowly (as expected under weak curvature) but converges to the same equilibrium law. This serves as a baseline check: the Fisher–Lyapunov equilibrium is quantitatively predictive, not merely qualitative. We observed early entry into the basin of Fisher geometry even from random starts; see Section 9.3 for a diagnostic based on the stability ratios $\delta_{F}$ and $\delta_{H}$.
Figure 2: OU benchmark: stationary variances vs. batch size $b$ (log scales). Solid: simulation; dashed: Lyapunov prediction from (48).
$d=2$, $H=\mathrm{diag}(1,0.1)$, $F=I_{2}$, 500 replicates, 200k steps.
8.1.2 Experiment 2: Anisotropic Fisher–Hessian Misalignment (Angle Sweep).
We keep the same curvature $H=\mathrm{diag}(1,0.1)$ but rotate an anisotropic noise-shape matrix relative to the curvature basis:
$F(\varphi)=R(\varphi)\begin{pmatrix}1.5&0\\[2.0pt] 0&0.5\end{pmatrix}R(\varphi)^{\top},\qquad\varphi\in[0,\pi/3],$
with fixed batch size, isolating alignment as an operational parameter. Because the simulator is Euler–Maruyama with finite time step, the appropriate comparator is the discrete Lyapunov equation: writing $A=I-\eta H\Delta t$ and $Q=(2\eta/b)F(\theta)\Delta t$, the stationary covariance solves
$\Sigma\;=\;A\Sigma A^{\top}+Q.$ (49)
Figure 3 shows a clear transfer of stationary variance across coordinates as $\varphi$ varies: when the noise eigenbasis rotates away from the curvature eigenbasis, equilibrium variance is redistributed in a predictable way, and the discrete Lyapunov benchmark tracks the simulation closely. The takeaway is structural: alignment governs which directions receive noise energy at equilibrium, even when scalar temperature $\eta/b$ is held fixed.
Figure 3: Angle sweep: stationary variances vs. misalignment angle. Bands show replicate variability; dashed curves are the discrete Lyapunov prediction from (49).
$d=2$, $H=\mathrm{diag}(1,0.1)$, 500 replicates per angle.
8.1.3 Experiment 4: OU/Lyapunov Plateau in $d=10$ (Anisotropic vs. Isotropic).
Experiments 1–2 used $d=2$. We now move to $d=10$ to validate the Lyapunov equilibrium in a setting where anisotropy is richer. The curvature is $H=\mathrm{diag}(\lambda_{1},\ldots,\lambda_{10})$ with eigenvalues log-spaced between 1 and 100; the noise shape $G^{\star}$ is a random SPD matrix with eigenvalues in $[0.5,5]$, and we compare an isotropic control $G_{\mathrm{iso}}=\sigma^{2}I$ with $\sigma^{2}=\mathrm{Tr}(G^{\star})/d$ (trace-matched). Under constant step size $\eta=0.5/\lambda_{\max}(H)$ and batch size $b=64$, we run R=2, [^0] R=2{,}000 replicates and record the Fisher-metric risk $\mathbb{E}[\|\theta_{T}\|_{F^{\star}}^{2}]$ over increasing horizons $T\in\{300,800,2000,5000,10000\}$.
Table 4 confirms the Lyapunov prediction (Corollary 5.7): both the anisotropic and isotropic empirical risks converge to the predicted stationary level as $T$ grows, with convergence ratios $\widehat{R}/R_{\infty}$ approaching unity from both sides. The two plateaux nearly coincide because trace-matching is exact in this diagonal-$H$ setting; directional differences emerge only in higher-order structure (Claim 3 below).
Empirical Risk $\widehat{R}_{T}$ Lyapunov $R_{\infty}$ $\widehat{R}_{T}/R_{\infty}$
$T$ Aniso Iso Aniso Iso Aniso Iso
300 $8.55\!\times\!10^{-4}$ $8.60\!\times\!10^{-4}$ $8.63\!\times\!10^{-4}$ $8.58\!\times\!10^{-4}$ 0.991 1.002
800 $8.64\!\times\!10^{-4}$ $8.49\!\times\!10^{-4}$ 1.002 0.990
2 000 $8.58\!\times\!10^{-4}$ $8.50\!\times\!10^{-4}$ 0.995 0.992
5 000 $8.72\!\times\!10^{-4}$ $8.58\!\times\!10^{-4}$ 1.010 1.001
10 000 $8.63\!\times\!10^{-4}$ $8.47\!\times\!10^{-4}$ 1.000 0.987
Table 4: OU/Lyapunov plateau ($d=10$): empirical Fisher-metric risk vs. discrete Lyapunov prediction $R_{\infty}=\mathrm{Tr}(F^{\star}\Sigma_{\infty})$. Convergence ratios $\widehat{R}_{T}/R_{\infty}\to 1$ confirm quantitative accuracy.
$d=10$, $H=\mathrm{diag}(1,\ldots,100)$, random $G^{\star}$ with eigenvalues in $[0.5,5]$, $\eta=0.5/\lambda_{\max}$, $b=64$, R=2, [^0] R=2{,}000 replicates.
8.2 Claim 2: Decaying Step Sizes Recover a $1/N$ Rate with a Geometry-Determined Constant
The second claim is that under a Robbins–Monro step-size schedule the Fisher-metric risk decays as $C/N$ where $N=Tb$ is the total oracle budget and the constant $C$ is determined by the intrinsic geometry $(G^{\star},H^{\star},F^{\star})$. Specifically, the SA central-limit theorem gives an asymptotic covariance $V$ satisfying $({\eta_{0}H^{\star}-\tfrac{1}{2}I})V+V({\eta_{0}H^{\star}-\tfrac{1}{2}I})=\frac{\eta_{0}^{2}}{b}\,G^{\star}$, and $N\cdot\mathrm{Risk}\to b\,\mathrm{Tr}(F^{\star}V)$. This validates the matching upper and lower bounds (Theorem 6.5 and Proposition 6.10).
8.2.1 Experiment 5: $1/N$ Rate and Fisher Constant (Decaying Step Size).
We switch to a decaying step size $\eta_{t}=\eta_{0}/(t+t_{0})$ with $\eta_{0}=1.2$, $t_{0}=50$, and the same $(H,G^{\star},G_{\mathrm{iso}})$ from Experiment 4, with oracle budget $N=Tb$ and $b=64$. Figure 4 (left) plots the Fisher-metric risk on log–log axes: both anisotropic and isotropic risks decay as $1/N$ (fitted slopes $\approx-0.98$), confirming Theorem 6.5. The right panel shows the scaled risk $N\cdot\widehat{\mathrm{Risk}}$, which stabilizes near $\approx 14.2$ (anisotropic) and $\approx 14.0$ (isotropic)—both close to the SA–CLT prediction $b\,\mathrm{Tr}(F^{\star}V)$ where $V$ solves the Lyapunov equation above. The inset marks the auxiliary constant $\mathrm{Tr}(G^{\star}H^{-1})\approx 5.41$, which governs the constant-step OU stationary risk (Corollary 5.7); the ratio of the decaying-step plateau to this constant reflects the step-size schedule and the interplay between curvature and noise geometry.
Figure 4: Decaying step size, $d=10$. Left: Fisher-metric risk vs. NN (log–log); slopes $\approx-1$ confirm $1/N$ rate. Right: Scaled risk $N\cdot\widehat{\mathrm{Risk}}$ stabilizes; inset shows $C=\mathrm{Tr}(G^{\star}H^{-1})$.
$\eta_{t}=1.2/(t+50)$, $b=64$, 800 replicates.
Table 5 summarizes the tightness of the bounds. The empirical risk is bounded between the information-theoretic lower bound (Proposition 6.10) and the upper bound (Theorem 6.5), with $N\cdot\text{Risk}$ stabilizing near $14.2$ (anisotropic) and $14.0$ (isotropic), both above the van Trees lower bound cd=10c\,d=10. At the largest budget (N=614,400N=614{,}400) the ratio of empirical constant to lower bound is approximately $1.4\times$, indicating near-tightness. The rate and constant are thus determined by $(G^{\star},H^{\star})$, so the oracle budget for a target risk can be computed from the problem’s intrinsic geometry.
Table 5: Upper vs. lower bound: scaled risk $N\cdot\widehat{\mathrm{Risk}}$ compared with the van Trees lower bound $cd=10$.
NN $N\!\cdot\!\widehat{\mathrm{Risk}}_{\mathrm{aniso}}$ $N\!\cdot\!\widehat{\mathrm{Risk}}_{\mathrm{iso}}$ $d=10$ (lower bound)
38 400 13.59 13.41 10
76 800 13.66 13.44 10
153 600 13.66 14.26 10
307 200 14.07 14.25 10
614 400 14.35 14.42 10
The SA–CLT Lyapunov prediction gives $b\,\mathrm{Tr}(F^{\star}V)\approx 14.2$ (anisotropic) and $\approx 14.0$ (isotropic); $N\cdot\mathrm{Risk}$ stabilizes above the van Trees lower bound cd=10c\,d=10 as expected.
8.3 Claim 3: Scalar Temperature Matching Fails—Directionality Is the Signature of Geometry
The third claim is that matching the total noise power (scalar temperature $\tau=\eta/b$) is insufficient: the directional distribution of noise energy is a distinguishing signature of the identified geometric model that no isotropic surrogate can reproduce. The Lyapunov trace identity $\mathrm{Tr}(H^{\star}\Sigma_{\infty})=\frac{\tau}{2}\mathrm{Tr}(G^{\star})$ depends only on $\mathrm{Tr}(G^{\star})$, so a trace-matched isotropic surrogate reproduces total risk by construction; but the directional distribution is entirely wrong, spreading residual error uniformly rather than concentrating it along top noise eigendirections.
8.3.1 Experiment 3: Geometry vs. Scalar Temperature (Rotation Diagnostic).
We compare a geometric OU model (diffusion covariance $\propto F_{\star}(\phi)$, rotating with angle $\phi$) against a trace-matched isotropic surrogate ($\sigma^{2}I$ with $\sigma^{2}=\mathrm{Tr}(F_{\star})/d$). At stationarity, the geometric covariance $\Sigma_{F}(\phi)$ generically rotates (producing nontrivial cross-covariance), while the isotropic $\Sigma_{I}(\phi)$ remains axis-aligned. Figures 57 confirm this: the ellipses tilt, cross-covariance $\Sigma_{12}$ tracks theory, and marginal variances redistribute—effects that temperature matching alone cannot capture.
Figure 5: Stationary covariance ellipses at three rotation angles: the identified equilibrium tilts with $\varphi$; the isotropic trace-matched surrogate remains axis-aligned.
$d=2$, $b=64$. Ellipses: 95% level sets of the Lyapunov solution.
Figure 6: Cross-covariance $\Sigma_{12}$ vs. rotation angle $\varphi$: isotropic trace-matched noise fails to reproduce off-diagonal covariance. Bands show $\pm 1$ s.d. across seeds.
$d=2$, 6 seeds, T=2,500T=2{,}500, burn-in=800{}=800.
Figure 7: Marginal variances $\Sigma_{11}$ (left) and $\Sigma_{22}$ (right) vs. angle: total noise energy may match, but directional allocation differs.
8.3.2 Experiment 6: Directional Variance Amplification.
We now move to $d=10$ and ask whether anisotropy affects how risk is distributed across directions. For $k=1,2,3$ we project $\theta_{T}$ onto the top eigenvectors $v_{k}$ of $G^{\star}$ and compute the directional second moment $\mathbb{E}[\langle v_{k},\theta_{T}\rangle^{2}]$. Figure 8 (left) displays these on log–log axes: both models decay as $1/N$, but the anisotropic curve is systematically above the isotropic one along the top eigendirections. The right panel plots the ratio $\mathbb{E}_{\mathrm{aniso}}/\mathbb{E}_{\mathrm{iso}}$; at the largest budget the ratios are $1.65$, $1.45$, and $1.26$ for $k=1,2,3$, mirroring the eigenvalue ordering $\lambda_{1}>\lambda_{2}>\lambda_{3}$ of $G^{\star}$. These amplification ratios show that the identified geometry determines which parameter directions carry the most residual risk—information that scalar temperature models discard.
Figure 8: Directional variance amplification ($d=10$). Left: Directional second moments along top eigenvectors of $G^{\star}$. Right: Ratio of anisotropic to isotropic directional variances; ratios >1>1 mirror the eigenvalue ordering of $G^{\star}$. Temperature matching preserves $\mathrm{Tr}(G^{\star})$ but not the eigen-structure; anisotropic noise concentrates residual risk along the top noise eigendirections.
Same as Experiment 5; top $k=1,2,3$ eigenvectors of $G^{\star}$.
Additional robustness experiments under misspecification and approximate exchangeability are reported in Appendix B.
9 Discussion
9.1 Summary of Contributions
The paper’s central contribution is the identification of the mini-batch gradient covariance from sampling design and exchangeability as the projected gradient-noise matrix $G^{\star}(\theta)$—which reduces to Fisher information at $\theta^{\star}$ under correct specification and, for general losses, induces Godambe/sandwich geometry together with $H^{\star}$. This identification endogenizes the diffusion coefficient, determines the Lyapunov stationary covariance, and selects the natural metric for rates and complexity, yielding matching $1/N$ bounds governed by intrinsic dimension $d_{\mathrm{eff}}$ and condition number $\kappa_{F}$.
  • Proven: Alignment (Theorem 4.3), diffusion (Theorem 5.4), OU/Lyapunov (Corollary 5.7), discrete bridge (Proposition 5.9), matching bounds (Theorem 6.5 and Proposition 6.10), and oracle complexity (Theorem 7.3).
  • Sketched: Batch-size scaling (Corollary 7.8) and variance-reduction bridge (Remark 9.1).
  • Deferred: Misspecification, weak dependence, explicit batching policies, and OR-specific applications (Appendix B provides preliminary experiments).
9.2 Implications for Algorithm Design
In the diffusion regime the key knob is the effective temperature $\tau=\eta/b$. Changing $b$ does more than “reduce variance”; it changes the diffusion amplitude in a directional way determined by the identified geometry, and therefore changes the Lyapunov-predicted error floor. Because the intrinsic noise is matrix-valued, variance-reduction interventions should be evaluated by their effect on risk in a chosen metric (e.g., $\mathrm{Tr}(Q\Sigma)$) rather than Euclidean variance alone, connecting classical OR variance-reduction practice to SGD through the correct geometry.
Remark 9.1 (Variance reduction as Fisher-metric risk reduction).
A VR technique that reduces $G^{\star}$ to $\widetilde{G}\preceq G^{\star}$ in the Loewner order yields $\widetilde{\Sigma}_{\infty}\preceq\Sigma_{\infty}$ via the Lyapunov balance, and hence lower Fisher-metric risk.
SGD does not precondition the drift; natural-gradient methods do so by reweighting $\nabla L$. Our results explain why plain SGD nevertheless exhibits geometry-aware behavior through its noise structure: drift contraction and diffusion injection are distinct levers. The control view also suggests adaptive batching—increasing $b$ (cooling) when marginal variance reduction is valuable in the performance metric of interest, and decreasing it when exploration matters.
9.3 Practical Notes and Extensions
[…]
Our analysis freezes $F(\theta)$ locally at $F^{\star}$. Validity can be assessed by monitoring $\delta_{F}(t):=\|F(\theta_{t})-F^{\star}\|_{\operatorname{op}}/\|F^{\star}\|_{\operatorname{op}}$; in our experiments the Lyapunov predictions remain accurate from random starts, indicating early entry into the basin of Fisher geometry.
9.4 Estimating $G^{\star}$ in Practice
The preceding theory identifies $G^{\star}(\theta)$ as the central primitive. In applications where $G^{\star}$ is not available analytically, several estimation strategies provide operationally useful approximations. The simplest is the empirical covariance of per-sample gradients: given a pilot batch $\{z_{1},\ldots,z_{m}\}$ at $\theta_{t}$, the sample covariance $\hat{G}(\theta_{t}):=\frac{1}{m-1}\sum_{i=1}^{m}(\psi_{i}-\bar{\psi})(\psi_{i}-\bar{\psi})^{\top}$ (where $\psi_{i}:=\nabla_{\theta}\ell(\theta_{t};z_{i})$) is the natural plug-in estimator and converges at the standard $O(1/\sqrt{m})$ rate in operator norm. For large $d$, a diagonal approximation $\hat{G}_{\mathrm{diag}}$ retains the per-coordinate noise scales and suffices for estimating $d_{\mathrm{eff}}$ and detecting gross anisotropy; when the leading eigendirections of $G^{\star}$ are the primary target, a rank-$k$ approximation via streaming PCA on per-sample gradients requires only $O(dk)$ storage. If $G^{\star}(\theta)$ varies slowly along the optimization path, an exponential moving average $\hat{G}_{t}=(1-\alpha)\hat{G}_{t-1}+\alpha\,(\psi_{t}-\bar{\psi}_{t})(\psi_{t}-\bar{\psi}_{t})^{\top}$ tracks the local geometry with minimal overhead.
[…]
9.5 Limitations and Future Work
Our sharpest statements use local linearization near a nondegenerate critical point; extending to nonconvex regimes requires controlling state-dependent curvature and relating intrinsic geometry to metastability. When $d\gg n$, the identified covariance may be low-rank, and a satisfactory theory should work on identifiable manifolds with appropriate pseudoinverses. Finally, preconditioners (Gauss–Newton, KFAC) reshape both drift and noise geometry; a joint control theory co-designing contraction and diffusion—potentially under communication constraints—would connect information geometry to OR-style variance allocation at scale.
Appendix A Technical Lemmas and Concentration Tools
This appendix collects the auxiliary lemmas invoked by the main-text proofs. All spine results—the alignment theorem, the diffusion limit, the OU linearization, the Lyapunov corollary, the upper and lower bounds, and the oracle complexity guarantee—are proved in the body of the paper.
A.1 Finite-population covariance identity
The following identity, classical in survey sampling (Cochran, 1977; Serfling, 1980), is used in the proof of the alignment theorem (Theorem 4.3).
Lemma A.1 (Without-replacement covariance of a mini-batch mean).
Let $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ be deterministic vectors with mean $\bar{x}_{n}:=\frac{1}{n}\sum_{i=1}^{n}x_{i}$ and centered sample covariance
$S_{n}:=\frac{1}{n-1}\sum_{i=1}^{n}(x_{i}-\bar{x}_{n})(x_{i}-\bar{x}_{n})^{\top}.$
Let $B$ be a subset of $\{1,\dots,n\}$ of size $b$ drawn uniformly without replacement, and define $\bar{x}_{B}:=\frac{1}{b}\sum_{i\in B}x_{i}$. Then, conditional on $(x_{1},\dots,x_{n})$,
$\mathrm{Cov}(\bar{x}_{B}\mid x_{1:n})=\frac{1}{b}\left(1-\frac{b}{n}\right)S_{n}.$ (50)
Proof.
Write $\bar{x}_{B}=\frac{1}{b}\sum_{i=1}^{n}I_{i}x_{i}$ where $I_{i}=\mathbf{1}\{i\in B\}$. Under SRSWOR, $\mathbb{E}[I_{i}]=b/n$, $\mathrm{Var}(I_{i})=\frac{b}{n}(1-\frac{b}{n})$, and $\mathrm{Cov}(I_{i},I_{j})=-\frac{b(n-b)}{n^{2}(n-1)}$ for $i\neq j$. Expanding $\mathrm{Cov}(\bar{x}_{B}\mid x_{1:n})=b^{-2}\sum_{i,j}\mathrm{Cov}(I_{i},I_{j})x_{i}x_{j}^{\top}$ and collecting terms:
$b^{-2}\!\left[\frac{b(n-b)}{n^{2}}\sum_{i}x_{i}x_{i}^{\top}-\frac{b(n-b)}{n^{2}(n-1)}\sum_{i\neq j}x_{i}x_{j}^{\top}\right]$
$\quad=\frac{n-b}{bn^{2}}\cdot\frac{n}{n-1}\sum_{i=1}^{n}(x_{i}-\bar{x}_{n})(x_{i}-\bar{x}_{n})^{\top}=\frac{1}{b}\!\left(1-\frac{b}{n}\right)S_{n},$
where the last equality uses $(n-1)S_{n}=\sum(x_{i}-\bar{x}_{n})(x_{i}-\bar{x}_{n})^{\top}$. ∎
A.2 Score identities and information equality
Lemma A.2 (Score identities and information equality).
Assume $\theta\mapsto p_{\theta}(z)$ is differentiable, $\int p_{\theta}(z)\,dz=1$, and differentiation under the integral sign is valid. Then:
  1. (i)
    $\mathbb{E}_{\theta}[s_{\theta}(Z)]=0$, where $s_{\theta}(z):=\nabla_{\theta}\log p_{\theta}(z)$ is the score.
  2. (ii)
    $F(\theta)=\mathbb{E}_{\theta}[s_{\theta}(Z)s_{\theta}(Z)^{\top}]=-\mathbb{E}_{\theta}[\nabla_{\theta}^{2}\log p_{\theta}(Z)]$ (information equality).
Proof.
Standard; see, e.g., Lehmann and Casella (1998, Ch. 7). ∎
A.3 Lyapunov equation solution
Lemma A.3 (Lyapunov equation solution).
Let $A\in\mathbb{R}^{d\times d}$ with all eigenvalues having strictly positive real part. Let $Q\succeq 0$. Then the continuous Lyapunov equation $A\Sigma+\Sigma A^{\top}=Q$ has the unique solution
$\Sigma=\int_{0}^{\infty}e^{-As}Qe^{-A^{\top}s}ds.$
Proof.
Convergence: Since all eigenvalues of $A$ have positive real part, $\|e^{-As}\|\leq Ce^{-\lambda s}$ for some $C,\lambda>0$, ensuring convergence.
Verification: $A\Sigma+\Sigma A^{\top}=\int_{0}^{\infty}\frac{d}{ds}(-e^{-As}Qe^{-A^{\top}s})ds=[-e^{-As}Qe^{-A^{\top}s}]_{0}^{\infty}=Q$.
Uniqueness: If $\Sigma_{1},\Sigma_{2}$ both solve the equation, then $\Delta:=\Sigma_{1}-\Sigma_{2}$ satisfies $A\Delta+\Delta A^{\top}=0$, whose only solution is $\Delta=0$ when $A$ has eigenvalues with positive real part. ∎
A.4 Product bound for Robbins–Monro weights
Lemma A.4 (Product bound for Robbins-Monro weights).
Let $a>1$ and $c_{0}>0$. Define
$\alpha_{k}:=1-\frac{a}{k}+\frac{c_{0}}{k^{2}},\qquad\Pi_{s,T}:=\prod_{k=s}^{T-1}\alpha_{k}.$
Then for all $T$ sufficiently large:
  1. (i)
    $\Pi_{1,T}\leq CT^{-a}$ for a constant $C=C(a,c_{0})$.
  2. (ii)
    $\Pi_{t+1,T}\leq C(t/T)^{a}$ for $1\leq t<T$.
Proof.
Taking logarithms: $\log\alpha_{k}=\log(1-a/k+c_{0}/k^{2})=-a/k+O(1/k^{2})$ for large $k$. Thus $\log\Pi_{s,T}=-a\sum_{k=s}^{T-1}k^{-1}+O(1)=-a(\log T-\log s)+O(1)$. Exponentiating: $\Pi_{s,T}=O((s/T)^{a})$. For $s=1$: $\Pi_{1,T}=O(T^{-a})$. For $s=t+1$: $\Pi_{t+1,T}=O((t/T)^{a})$. ∎
A.5 Freedman’s inequality
Lemma A.5 (Freedman’s inequality).
Let $(M_{t})_{t\geq 0}$ be a martingale with $M_{0}=0$ and increments $\Delta M_{t}:=M_{t}-M_{t-1}$ satisfying $|\Delta M_{t}|\leq R$ almost surely. Let $V_{T}:=\sum_{t=1}^{T}\mathbb{E}[(\Delta M_{t})^{2}\mid\mathcal{F}_{t-1}]$ be the predictable quadratic variation. Then for any $v>0$ and $\lambda>0$:
$\Pr(M_{T}\geq\lambda\text{ and }V_{T}\leq v)\leq\exp\left(-\frac{\lambda^{2}}{2v+2R\lambda/3}\right).$
Equivalently, with probability at least $1-\delta$, whenever $V_{T}\leq v$:
$M_{T}\leq\sqrt{2v\log(1/\delta)}+\frac{2R}{3}\log(1/\delta).$
Proof.
See Freedman (1975) or Tropp (2015, Theorem 1.6). ∎
Appendix B Robustness Beyond Pure Exchangeability
The experiments in this appendix are exploratory robustness diagnostics, not part of the formal theoretical validation in Section 8. They demonstrate that the covariance-identification theory degrades gracefully under realistic departures from i.i.d. likelihood assumptions. Experiment A considers model misspecification, where curvature and noise geometry diverge; Experiment B models mild violations of exchangeability via variance inflation.
B.1 Experiment A: Misspecification (Probit DGP, Logistic Fit)
We generate n=50,000n=50{,}000 observations from a probit DGP and fit a logistic model, so that the optimization target is the KL projection $\beta^{\star}$ rather than the true $\beta_{0}$. Under misspecification the score covariance and curvature no longer coincide: $J^{\star}:=\mathrm{Var}\!\bigl(\nabla_{\beta}\ell(\beta;X,Y)\big|_{\beta=\beta^{\star}}\bigr)\neq H^{\star}:=\nabla^{2}\mathbb{E}[-\ell(\beta;X,Y)]\big|_{\beta=\beta^{\star}}$. The covariance-identification theorem (Theorem 4.3) nevertheless predicts $\mathrm{Cov}(g_{B}(\beta^{\star}))\approx(1/b)\,J^{\star}$, where $J^{\star}$ plays the role of $G^{\star}(\theta^{\star})$ in the general-loss notation. We estimate $J^{\star}$ from the full-sample scores and compare its diagonal entries to those of the empirical mini-batch covariance (batch size $b=256$, 800 replicates). Figure 9 shows close agreement (relative Frobenius difference $\approx 0.05$), confirming that the identification remains accurate under misspecification: mini-batch fluctuations are governed by the score covariance $J^{\star}$ rather than by curvature alone.
Figure 9: Experiment A: misspecification (probit DGP, logistic fit). Diagonal entries of the empirical mini-batch covariance vs. $(1/b)\,\hat{J}^{\star}$ (projected score covariance at $\beta^{\star}$); relative Frobenius difference $\approx 0.05$.
B.2 Experiment B: Approximate Exchangeability and Variance Inflation
To model mild departures from ideal exchangeability (e.g., weak dependence, contamination, or partial reuse of samples), we inflate the intrinsic covariance by a factor $(1+\varepsilon)$ in the linearized surrogate: $\mathrm{Cov}(\xi_{t})=(1/b)(1+\varepsilon)\,G^{*}(\theta^{\star})$, $\varepsilon\in[0,1]$. This captures “extra correlation” as increased effective diffusion amplitude. Linearity of the Lyapunov operator then implies $\Sigma(\varepsilon)=(1+\varepsilon)\Sigma(0)$, so every stationary marginal variance inflates linearly in $\varepsilon$. Figure 10 confirms this prediction: the entire stationary covariance scales as though the system had a higher effective temperature $\tau_{\mathrm{eff}}=(1+\varepsilon)\eta/b$. To first order, small exchangeability violations act like “heating” the Fisher–Lyapunov equilibrium.
Figure 10: Experiment B: approximate exchangeability. Stationary variances vs. $\varepsilon$ (simulation bands vs. linear Lyapunov prediction).
Appendix C Reproducibility and Code Availability
All experiments are reproducible from the accompanying code archive. The Python suite (Python_Codes/) includes a master script run_all.py that executes all experiments sequentially. Every script uses explicit random seeds (global seed =7=7) for deterministic reproducibility; small platform-level floating-point differences may produce minor variations ($\lesssim 1\%$) in Monte Carlo estimates.
Software environment.
Python $\geq$3.9; NumPy, SciPy, Matplotlib, Pandas. Tested under Python 3.10.12, NumPy 2.2, SciPy 1.15, Matplotlib 3.10. Total wall-clock time: $\approx$4 minutes on a single CPU core.
Script inventory.
Table 6 maps each script to the figures and tables it generates.
Table 6: Code inventory: scripts, outputs, and paper cross-references.
Script Experiment Figures / Tables Sec.
partA_batch_scaling.py 1 (Batch scaling) Fig. 2 8.1
partB_anisotropic_alignment.py 2 (Angle sweep) Fig. 3 8.1
partC_rotation_sweep.py 3 (Rotation) Figs. 57 8.3
bounds.py 4–6 (OU plateau, rate, directional) Tab. 4, Figs. 48, Tab. 5 8.18.3
partD_misspecification.py A (Misspecification) Fig. 9 App. B
partE_eps_exchangeable.py B ($\varepsilon$-exchangeable) Fig. 10 App. B
The script run_all.py executes all of the above in sequence. The script partF_optimal_batching.py produces additional exploratory figures not included in the paper.
Shared parameters ($d=10$ experiments).
Hessian eigenvalues log-spaced in $[1,100]$; $F^{\star}=H^{\star}$ (quadratic Gaussian model). Anisotropic gradient covariance $G^{\star}$ drawn as a random SPD matrix with eigenvalues in $[0.5,5]$ (seed =7=7); isotropic baseline $G_{\mathrm{iso}}=\sigma^{2}I$ with $\sigma^{2}=\mathrm{Tr}(G^{\star})/d$ (trace-matched). Constant-step experiments use $\eta=0.5/\lambda_{\max}(H)$, $b=64$, $\tau=\eta/b$; decaying-step experiments use $\eta_{t}=1.2/(t+50)$, $b=64$.
Verification.
All scripts were executed to completion with zero errors. Numerical outputs confirm: (i) log-log slopes $\approx-1$ for the $1/N$ risk decay; (ii) OU stationary variances match Lyapunov predictions to $1\%$; (iii) directional variance ratios align with the eigenstructure of $G^{\star}$; (iv) misspecification experiments show graceful degradation to the sandwich/Godambe prediction.